### v1

# A Novel Testing Approach for Differences Among Brain Connectomes

Abstract.
Statistical analysis on non-Euclidean spaces typically relies on distances as the primary tool for constructing likelihoods. However, manifold-valued data admits richer structures in addition to Riemannian distances. We demonstrate that simple, tractable models that do not rely exclusively on distances can be constructed on the manifold of symmetric positive definite (SPD) matrices, which naturally arises in brain connectivity analysis. Specifically, we highlight the manifold-valued Mahalanobis distribution, a parametric family that extends classical multivariate concepts to the SPD manifold. We develop estimators for this distribution and establish their asymptotic properties. Building on this framework, we propose a novel ANOVA test that leverages the manifold structure to obtain a test statistic that better captures the dimensionality of the data. We theoretically demonstrate that our test achieves superior statistical power compared to distance-based Fréchet ANOVA methods.
Key words and phrases:
ANOVA, Connectomics, Riemannian geometry
1. Introduction
The analysis of data that takes values on non-Euclidean spaces has emerged as a fundamental challenge across numerous scientific disciplines, from neuroimaging to computer vision. This has driven the development of sophisticated statistical tools specifically designed for data taking values on manifolds and metric spaces.
The statistical analysis of manifold-valued data now encompasses a rich toolkit. Fréchet ANOVA (muller) extends classical analysis of variance to general metric spaces, while functional data analysis has been adapted to handle curves and shapes on manifolds (fletcher; cornea). Shape analysis methods provide powerful frameworks for comparing anatomical structures (zhao), and Riemannian techniques have proven particularly effective for analyzing brain connectivity data (logeu; LRDF06; varo).
While metric space methods offer broad generality, real-world data often possess richer structure than just distances. Many datasets naturally lie on smooth manifolds equipped with differentiable structure: brain connectivity matrices live in symmetric positive definite (SPD) manifolds, directional data lie on spheres, and shape spaces have well-understood Riemannian geometry. By exploiting this additional structure, we can develop more powerful statistical methods tailored to the specific geometry of the data.
This specialization represents a deliberate trade-off: we sacrifice the universality of metric space approaches to gain statistical power through geometric tools. Methods that leverage manifold structure can exploit tangent space approximations, parallel transport, and geodesic regression, which are absent in general metric spaces. While this may require computing gradients and other differential quantities, modern computational resources and automatic differentiation make these calculations tractable even for large-scale problems.
In this paper, we focus on developing such specialized methods for brain connectivity analysis. Brain connectomes naturally form a Riemannian manifold with well-studied geometry. The Affine Invariant Riemannian Metric (AIRM) on the SPD manifold $\mathcal{M}$ provides a principled framework for analyzing these data (logeu; LRDF06).
Statistical inference on connectome-valued data has progressed significantly. The sample Fréchet mean has been shown to possess desirable properties including existence, uniqueness for negatively-curved manifolds (karcher; kendall), consistency (batta), and asymptotic normality (kendall2; clt1). Building on these foundations, researchers have developed regression methods for manifold-valued outcomes (fletcher; cornea) and comparative analyses across populations (varo).
However, existing methods for comparing groups of connectomes are understudied. The popular Fréchet ANOVA test (muller), while applicable to general metric spaces, relies on comparing variations (sums of squared distances from data points to group means). Critically, when applied to Euclidean data, Fréchet ANOVA does not reduce to classical MANOVA tests like Wilks’ Lambda or Pillai’s trace. We propose a Riemannian generalization of classical MANOVA. Our approach leverages the differentiable structure of the SPD manifold to construct test statistics that reduce to their classical counterparts in Euclidean space while remaining computationally tractable through modern optimization techniques.
Our contributions consist of three parts. First, we introduce a test statistic that generalizes classical Wilks’ Lambda to the SPD manifold, providing a principled bridge between Euclidean and Riemannian inference. Second, we derive the asymptotic distributions of these statistics. Third, we theoretically demonstrate that our approach achieves greater statistical power than existing methods under certain assumptions.
The remainder of this paper is organized as follows. In Section 2, we develop our methodological framework in three stages. We begin by introducing parameter estimation for the Mahalanobis Affine Invariant Riemannian Metric (MAIRM), a generalization of the classical affine invariant metric that relaxes sphericity assumptions. We then construct our Riemannian generalization of Wilks’ Lambda, establishing its fundamental properties and showing how it reduces to classical MANOVA in the Euclidean case. Finally, we derive the asymptotic distribution of our test statistic under the null hypothesis and analyze its behavior under alternatives, formally comparing its power properties to those of Fréchet ANOVA. Section 3 summarizes our contributions and discusses extensions for future work. Technical details on the Riemannian geometry of the SPD manifold and additional computational formulas are provided in the appendices.
2. Methods
2.1. Parameter Estimation
In a number of statistical models based on Riemannian geometry ideas, it is assumed that the likelihood depends exclusively on the geodesic distance between observations and some reference point. The simplest such model is decribed in varo: for a manifold-valued random variable $\boldsymbol{C}$,
$p(\boldsymbol{C})\propto\exp\left(-\frac{1}{2\sigma^{2}}\norm{\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C})}_{\boldsymbol{C}^{\ast}}^{2}\right)$
for some $\boldsymbol{C}^{\ast}\in\mathcal{M},\sigma\in\mathbb{R}^{+}$. Here, $\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C})$ denotes the Riemannian logarithm map. Inference for such a model is simple: given an iid sample $\{\boldsymbol{C}_{i}\}_{i}$, the MLE estimators for $\boldsymbol{C}^{\ast}$ and $\sigma^{2}$ are the sample Fréchet mean, $\bar{\boldsymbol{C}}_{n}$ and
$\hat{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}\norm{\log_{\bar{\boldsymbol{C}}_{n}}(\boldsymbol{C}_{i})}_{\bar{\boldsymbol{C}}_{n}}^{2}$
The estimators are independent.
Such a model makes a sphericity assumption. To make it explicit, consider the AIRM metric for concreteness and recall the vectorization maps (logeu):
$\operatorname{Vect}_{I}:T_{I}\mathcal{M}$ $\rightarrow\mathbb{R}^{p(p+1)/2}$
$\{w_{ij}\}$ $\mapsto(w_{11},\sqrt{2}w_{12},w_{22},\ldots)$
and
$\operatorname{Vect}_{\boldsymbol{C}}:T_{\boldsymbol{C}}\mathcal{M}$ $\rightarrow\mathbb{R}^{p(p+1)/2}$
$\log_{\boldsymbol{C}}(\boldsymbol{\Lambda})$ $\mapsto\operatorname{Vect}_{I}(\boldsymbol{C}^{-1/2}\log_{\boldsymbol{C}}(\boldsymbol{\Lambda})\boldsymbol{C}^{-1/2})$
Here we have used the identification of $T_{\boldsymbol{C}}\mathcal{M}$ with the vector space of $p$ by $p$ symmetric matrices. Both maps are linear isometries.
The model can be rewritten using these maps as
$p(\boldsymbol{C})\propto\exp\left(-\frac{1}{2\sigma^{2}}\operatorname{Vect}_{\boldsymbol{C}^{\ast}}(\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C}))^{T}\operatorname{Vect}_{\boldsymbol{C}^{\ast}}(\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C}))\right)$
In other words $\operatorname{Vect}_{\boldsymbol{C}^{\ast}}(\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C}))\sim N_{p(p+1)/2}(0,\sigma^{2}I)$.
Let us relax that assumption and consider the model
$p(\boldsymbol{C})\propto\exp\left(-\frac{1}{2}\operatorname{Vect}_{\boldsymbol{C}^{\ast}}(\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C}))^{T}\Gamma^{-1}\operatorname{Vect}_{\boldsymbol{C}^{\ast}}(\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C}))\right)$
for some positive definite $\Gamma$.
Given an iid sample $\{\boldsymbol{C}_{i}\}$, it is not hard to come up with natural estimators for $\boldsymbol{C}^{\ast}$ and $\Gamma$. First, notice that $\boldsymbol{C}^{\ast}$ is still the population Fréchet mean, so the sample Fréchet mean $\bar{\boldsymbol{C}}_{n}$ can still be used to estimate it. As for $\Gamma$, the estimator would be
$\hat{\mathbf{S}}_{n}=\frac{1}{n-1}\sum_{i=1}^{n}\operatorname{Vect}_{\bar{\boldsymbol{C}}_{n}}(\log_{\bar{\boldsymbol{C}}_{n}}(\boldsymbol{C}_{i}))\operatorname{Vect}_{\bar{\boldsymbol{C}}_{n}}(\log_{\bar{\boldsymbol{C}}_{n}}(\boldsymbol{C}_{i}))^{T}$
The price we pay for the added generality of this model is that it is no longer clear that $\bar{\boldsymbol{C}}_{n}$ and $(n-1)/n\hat{\mathbf{S}}_{n}$ are MLE estimators for $\boldsymbol{C}^{\ast}$ and $\Gamma$.
However, the fact that $\bar{\boldsymbol{C}}_{n}$ converges in probability to $\boldsymbol{C}^{\ast}$ is a direct consequence of Theorem 2.3 in clt4. In addition:
Proposition 1.
$\hat{\mathbf{S}}_{n}$ converges in probability to $\Gamma$.
Proof.
We have the chain of implications:
$\bar{\boldsymbol{C}}_{n}\overset{P}{\rightarrow}\boldsymbol{C}^{\ast}$
$\Rightarrow\bar{\boldsymbol{C}}_{n}^{-1/2}\boldsymbol{C}_{i}\bar{\boldsymbol{C}}_{n}^{-1/2}\overset{P}{\rightarrow}(\boldsymbol{C}^{\ast})^{-1/2}\boldsymbol{C}_{i}(\boldsymbol{C}^{\ast})^{-1/2}$
$\Rightarrow\log(\bar{\boldsymbol{C}}_{n}^{-1/2}\boldsymbol{C}_{i}\bar{\boldsymbol{C}}_{n}^{-1/2})\overset{P}{\rightarrow}$
$\qquad\log((\boldsymbol{C}^{\ast})^{-1/2}\boldsymbol{C}_{i}(\boldsymbol{C}^{\ast})^{-1/2})$
$\Rightarrow\operatorname{Vect}_{\bar{\boldsymbol{C}}_{n}}(\log_{\bar{\boldsymbol{C}}_{n}}(\boldsymbol{C}_{i}))\overset{P}{\rightarrow}\operatorname{Vect}_{\boldsymbol{C}^{\ast}}(\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C}_{i}))$
from which the result follows easily. ∎
A sphericity assumption is also used in regression models with Euclidean predictors and connectome-valued responses. A popular model (fletcher) is based on choosing $\boldsymbol{C}^{\ast}\in\mathcal{M}$, $\{v_{k}\}_{k}\in T_{\boldsymbol{C}^{\ast}}\mathcal{M}$, $\sigma^{2}\in\mathbb{R}^{+}$ and assumming that the observations $\{(\mathbf{x}_{i},\boldsymbol{C}_{i})\}_{i}$ are distributed according to the likelihood
$p(\boldsymbol{C}_{i})\propto\exp\left(-\frac{1}{2\sigma^{2}}\norm{\log_{\boldsymbol{C}_{i}^{\ast}}(\boldsymbol{C}_{i})}_{\boldsymbol{C}_{i}^{\ast}}^{2}\right)$
where $\boldsymbol{C}^{\ast}_{i}=\operatorname{Exp}_{\boldsymbol{C}^{\ast}}(\sum_{k}x_{k}v_{k})$.
Existence and uniqueness of OLS estimates in general is known (fletcher). Of course, with this likelihood they are MLE estimators.
We may relax the sphericity assumption by assumming instead that there is a positive definite $\Gamma$ such that
(1) $p(\boldsymbol{C}_{i})\propto\exp\left(-\frac{1}{2}\operatorname{Vect}_{\boldsymbol{C}_{i}^{\ast}}(\log_{\boldsymbol{C}_{i}^{\ast}}(\boldsymbol{C}_{i}))^{T}\boldsymbol{\Gamma}^{-1}\operatorname{Vect}_{\boldsymbol{C}_{i}^{\ast}}(\log_{\boldsymbol{C}_{i}^{\ast}}(\boldsymbol{C}_{i}))\right)$
In this case, we have
Proposition 2.
There exist unique MLE estimators for model (1)
Proof.
Consider the affine invariant Riemannian metric $\langle,\rangle^{\operatorname{Mah}}$, defined on $\mathcal{M}$ by requiring
$\langle\log_{\mathbf{I}}(\boldsymbol{C}),\log_{\mathbf{I}}(\boldsymbol{\Lambda})\rangle_{\mathbf{I}}^{\operatorname{Mah}}=\operatorname{Vect}_{\mathbf{I}}(\log_{\mathbf{I}}(\boldsymbol{C}))^{T}\boldsymbol{\Gamma}^{-1}\operatorname{Vect}_{\mathbf{I}}(\log_{\mathbf{I}}(\boldsymbol{\Lambda}))$
Then the MLE estimators of model (1) are the OLS estimates of $\langle,\rangle^{\operatorname{Mah}}$, whose existence and uniqueness follow from (fletcher). ∎
We call the family of metrics introduced in the previous proof Mahalanobis Affine Invariant Riemannian Metric (MAIRM).
The question of estimation rises again. The procedure we propose is similar to the previous one. The $\boldsymbol{C}_{i}^{\ast}$ still satisfy
$\boldsymbol{C}_{i}^{\ast}=\mathbb{E}[\boldsymbol{C}_{i}|\mathbf{X}=\mathbf{x}]$
hence the OLS predicted values $\hat{\boldsymbol{C}}_{i}$ are consistent estimators of $\boldsymbol{C}^{\ast}_{i}$ and
(2) $\hat{\mathbf{S}}_{n}=\frac{1}{n-1}\sum_{i}\operatorname{Vect}_{\hat{\boldsymbol{C}}_{i}}(\log_{\hat{\boldsymbol{C}}_{i}}(\boldsymbol{C}_{i}))\operatorname{Vect}_{\hat{\boldsymbol{C}}_{i}}(\log_{\hat{\boldsymbol{C}}_{i}}(\boldsymbol{C}_{i}))^{T}$
is a consistent estimator of $\Gamma$.
Thus, the MAIRM and AIRM predicted values are the same, but the confidence intervals are different.
2.2. Riemannian Generalization of Wilk’s Lambda
Let us now extend the concepts above to develop a multivariate test statistic that generalizes the classical Wilk’s Lambda to the SPD manifold. This generalization enables us to test for differences among multiple groups of connectomes while respecting the natural geometric structure of the data.
To introduce the statistic that generalizes Wilk’s Lambda, let us set up some notation. In the manifold $\mathcal{M}=\operatorname{SPD}$, consider $\mathcal{M}$-valued random variables $C_{lj}$, $l=1,\ldots,g$, $j=1,\ldots,n_{l}$.
The usual orthogonal decomposition that is the basis of ANOVA does not directly apply in this manifold setting, and we need to find a suitable substitute.
Let $\hat{C}$ denote the Fréchet sample mean of the whole sample and $\hat{C}_{l}$ denote the Fréchet sample mean of the subsample $\{C_{lj}\}_{j}$.
Proposition 3.
[…]
$T$ $=\sum_{l=1}^{g}\sum_{j=1}^{n_{l}}w_{lj}w_{lj}^{T}$
$W$ $=\sum_{l=1}^{g}\sum_{j=1}^{n_{l}}u_{lj}u_{lj}^{T}$
If $T$ is positive definite, then the test statistic
$\Lambda^{\ast}=\frac{|W|}{|T|}$
ranges between 0 and 1.
Proof.
Since the vectors $\{u_{lj}\}_{j}$ are centered (i.e., $\sum_{j=1}^{n_{l}}u_{lj}=0$), we have
$\sum_{j=1}^{n_{l}}w_{lj}w_{lj}^{T}=\sum_{j=1}^{n_{l}}(v_{l}-u_{lj})(v_{l}-u_{lj})^{T}$
(3) $=n_{l}v_{l}v_{l}^{T}+\sum_{j=1}^{n_{l}}u_{lj}u_{lj}^{T}$
Adding over $l$ yields the decomposition
(4) $T=\sum_{l=1}^{g}n_{l}v_{l}v_{l}^{T}+W$
Both matrices are positive semidefinite and equation (4) shows $T\geq W$, hence $0\leq\Lambda^{\ast}\leq 1$. ∎
Remark. Notice the difference between $w_{lj}$ and the vector $\operatorname{Vec}_{\hat{C}}(\log_{\hat{C}}C_{lj})$. In general, they are not the same, but in the Euclidean case they coincide, so the subsequent developments reduce to the traditional MANOVA in that case.
We consider the null hypothesis $H_{0}:C_{1}^{*}=C_{2}^{*}=\cdots=C_{g}^{*}$, where $C_{l}^{*}$ denotes the population Fréchet mean of group $l$. This is equivalent to testing whether all groups share the same population mean in the manifold.
2.3. Asymptotic Distribution of the Test Statistic
We now present the asymptotic distribution of our Riemannian ANOVA test statistic $\Lambda^{*}$ under this model.
Theorem 4 (Asymptotic Distribution).
Let $Y_{lj}\sim p(C)$ for $l=1,\ldots,g$ and $j=1,\ldots,n_{l}$, where
$p(C)\propto\exp\left(-\frac{1}{2}\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(C))^{T}\Gamma^{-1}\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(C))\right)$
is the Riemannian normal distribution on the manifold of symmetric positive definite matrices with population mean $C^{*}$ and covariance $\Gamma$ in the tangent space.
Under the null hypothesis $H_{0}:C^{*}_{1}=C^{*}_{2}=\cdots=C^{*}_{g}=C^{*}$ as $n=\sum_{l=1}^{g}n_{l}\to\infty$ with $n_{l}/n\to\pi_{l}\in(0,1)$:
$n\log\Lambda^{*}\xrightarrow{d}\chi^{2}_{d(g-1)}$
where $d=p(p+1)/2$ is the dimension of the tangent space for $p\times p$ symmetric positive definite matrices.
Proof.
The proof proceeds by establishing the asymptotic behavior of the components of $\Lambda^{*}$ and then deriving the limiting distribution. Recall from Proposition 3 that the test statistic is defined as:
$\Lambda^{*}=\frac{|W|}{|T|}$
where $T=\sum_{l=1}^{g}\sum_{j=1}^{n_{l}}w_{lj}w_{lj}^{T}$ and $W=\sum_{l=1}^{g}\sum_{j=1}^{n_{l}}u_{lj}u_{lj}^{T}$.
We begin by applying the central limit theorem for Fréchet means (clt4). Define the chart $\phi:\text{SPD}(p)\to\mathbb{R}^{d}$ by $\phi(C)=\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(C))$.
We first calculate the gradient $\Psi(u;\theta):=\text{grad}_{\theta}\rho_{\phi}^{2}(u,\theta)$, where $\rho_{\phi}^{2}(u,\theta)=\rho^{2}(\text{Exp}_{C^{*}}(u),\text{Exp}_{C^{*}}(\theta))$ and $\rho$ denotes the geodesic distance on the manifold. We have
$\Psi(u;\theta)=-2G_{C^{*}}(u-\theta)+O(\|\theta\|^{2})$
where $G_{C^{*}}$ is the matrix representation of the metric tensor at $C^{*}$ in vectorized coordinates, arising from the inner product $\langle V,W\rangle_{C^{*}}=\text{tr}((C^{*})^{-1}V(C^{*})^{-1}W)$.
Next, under $H_{0}$, with $u=\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(Y))\sim N(0,\Gamma)$, we have:
$\Sigma=\text{Cov}(\Psi(u;0))=\text{Cov}(-2G_{C^{*}}u)=4G_{C^{*}}\Gamma G_{C^{*}}^{T}$
We now compute the Hessian $\Lambda$. The Hessian of the squared geodesic distance at the minimum is:
$\Lambda_{ij}=E\left[\frac{\partial}{\partial\theta_{j}}\Psi_{i}(u;\theta)\Big|_{\theta=0}\right]=2[G_{C^{*}}]_{ij}$
That is, $\Lambda=2G_{C^{*}}$.
Applying Theorem 2.1 of batta, we obtain:
$\sqrt{n}(\phi(\hat{C})-\phi(C^{*}))\xrightarrow{d}N(0,\Lambda^{-1}\Sigma(\Lambda^{T})^{-1})$
Substituting our computed values:
$\Lambda^{-1}\Sigma(\Lambda^{T})^{-1}=(2G_{C^{*}})^{-1}\cdot 4G_{C^{*}}\Gamma G_{C^{*}}^{T}\cdot(2G_{C^{*}}^{T})^{-1}=\Gamma$
That is, the metric terms elegantly cancel, which establishes:
(5) $)\displaystyle\sqrt{n}\,\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(\hat{C}$)) $\xrightarrow{d}N(0,\Gamma)$
With this result at hand, we now turn to the asymptotic behavior of tangent vectors. For the between-group deviation vectors:
$)v_{l}=\operatorname{Vec}_{\hat{C}_{l}}(\log_{\hat{C}_{l}}(\hat{C}$))
we proceed as follows.
First, applying Corollary 9 with $A=\hat{C}_{l}$, $B=\hat{C}$, $C=C^{*}$, and noting that $d_{R}(\hat{C}_{l},C^{*})=O_{p}(n_{l}^{-1/2})$ and $d_{R}(\hat{C},C^{*})=O_{p}(n^{-1/2})$ from the convergence of Fréchet sample means, we obtain
$\log_{\hat{C}_{l}}(\hat{C})=\Pi_{C^{*}\to\hat{C}_{l}}[\log_{C^{*}}(\hat{C})-\log_{C^{*}}(\hat{C}_{l})]+O_{p}(n^{-1/2})$
By the central limit theorem for Fréchet means we have $\log_{C^{*}}(\hat{C})=O_{p}(n^{-1/2})$ and $\log_{C^{*}}(\hat{C}_{l})=O_{p}(n_{l}^{-1/2})$. Therefore, the difference becomes:
$\log_{C^{*}}(\hat{C})-\log_{C^{*}}(\hat{C}_{l})=O_{p}(n^{-1/2})$
Combining all these results:
$)\displaystyle v_{l}=\operatorname{Vec}_{\hat{C}_{l}}(\log_{\hat{C}_{l}}(\hat{C}$)) $=\operatorname{Vec}_{C^{*}}\Pi_{C^{*}\to\hat{C}_{l}}(\log_{C^{*}}(\hat{C})-\log_{C^{*}}(\hat{C}_{l}))+O_{p}(n^{-1/2})$
$=O_{P}(n^{-1/2})+O_{p}(n^{-1/2})=O_{P}(n^{-1/2})$
We now consider the observation vectors:
$u_{lj}=\operatorname{Vec}_{\hat{C}_{l}}(\log_{\hat{C}_{l}}(Y_{lj}))$
To establish the stability under change of base point, we analyze how the vectorization changes from $\hat{C}_{l}$ to $C^{*}$. Corollary 9 yields the expansion:
$\log_{\hat{C}_{l}}(Y_{lj})=\Pi_{C^{*}\to\hat{C}_{l}}\left(\log_{C^{*}}(Y_{lj})-\log_{C^{*}}(C_{l})\right)+O(n^{-1/2})$
Therefore, combining these estimates:
$u_{lj}=\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(Y_{lj}))+O_{p}(n_{l}^{-1/2})$
Next, we examine the decomposition and matrix distributions. The decomposition $T=B+W$ holds where:
(6) $B$ $=\sum_{l=1}^{g}n_{l}v_{l}v_{l}^{T}=O_{p}(1)$
(7) $W$ $=\sum_{l=1}^{g}\sum_{j=1}^{n_{l}}u_{lj}u_{lj}^{T}$
Due to the asymptotic normality in the tangent space, we have an approximate Wishart distribution for $W$. More precisely, the argument above gives:
$u_{lj}=z_{lj}+\varepsilon_{lj}$
where $z_{lj}\sim N(0,\Gamma)$ exactly and $\varepsilon_{lj}=O_{p}(n_{l}^{-1/2})$ represents the error from using the estimated group mean. This leads to the decomposition:
$W=\sum_{l=1}^{g}\sum_{j=1}^{n_{l}}u_{lj}u_{lj}^{T}=Z+E+R$
where $Z\sim\mathcal{W}_{d}(n-g,\Gamma)$ (the degrees of freedom are due to the centering constraints $\sum_{j=1}^{n_{l}}u_{lj}=0$) is the exact Wishart component, $E=O_{p}(n_{l}^{-1/2})$ represents cross-terms between true and error components, and $R=O_{p}(n_{l}^{-1/4})$ is the quadratic error term.
The rest of the argument proceeds exactly as in Euclidean MANOVA. ∎
2.4. Comparison to Fréchet ANOVA
The Fréchet ANOVA test (muller) provides a distribution-free alternative for testing equality of population Fréchet means and variances in general metric spaces. We now establish the precise relationship between Fréchet ANOVA and our Riemannian ANOVA test under the Riemannian normal model.
The Fréchet ANOVA test statistic takes the form:
$T_{n}=\frac{nU_{n}}{\sum_{l=1}^{g}\frac{\lambda_{l,n}}{\hat{\sigma}_{l}^{2}}}+\frac{nF_{n}^{2}}{\sum_{l=1}^{g}\lambda_{l,n}^{2}\hat{\sigma}_{l}^{2}}$
[…]
Proposition 5 (Fréchet ANOVA Non-centrality Under Normality).
Under the Riemannian normal model with equal covariances and alternative hypothesis $H_{1}$, the Fréchet ANOVA test statistic follows asymptotically:
$T_{n}\xrightarrow{d}\chi^{2}_{g-1}(\delta_{\text{Fréchet}})$
where the non-centrality parameter is:
$\delta_{\text{Fréchet}}=\frac{n\left(\sum_{l=1}^{g}\lambda_{l}d^{2}(C_{l}^{*},C^{*})\right)^{2}}{\operatorname{tr}(\Gamma)\sum_{l=1}^{g}\lambda_{l}^{2}}$
with $C^{*}$ denoting the overall population Fréchet mean.
Proof.
Suppose $Y_{lj}\sim p(C_{l}^{*})$ where the connectomes follow a Riemannian normal distribution on $\mathrm{SPD}(p)$ with population mean $C_{l}^{*}$ and covariance $\Gamma$ in the tangent space (as defined in Section 2.3).
Under the equal covariance assumption, the population Fréchet variance for each group is:
$V_{l}=\mathbb{E}[d^{2}(Y_{lj},C_{l}^{*})]=\operatorname{tr}(\Gamma)$
for all $l=1,\ldots,g$, since $\log_{C_{l}^{*}}(Y_{lj})\sim\mathcal{N}(0,\Gamma)$ in the vectorized tangent space.
The overall population Fréchet variance admits the decomposition:
$V_{p}=\sum_{l=1}^{g}\lambda_{l}\mathbb{E}[d^{2}(Y_{lj},C^{*})]=\sum_{l=1}^{g}\lambda_{l}\left(V_{l}+d^{2}(C_{l}^{*},C^{*})\right),$
(similar to equation (3)). This yields:
$V_{p}=\operatorname{tr}(\Gamma)+\sum_{l=1}^{g}\lambda_{l}d^{2}(C_{l}^{*},C^{*}).$
Consequently, the scale component of the Fréchet ANOVA statistic vanishes:
$U=\sum_{j<l}\frac{\lambda_{j}\lambda_{l}}{\sigma_{j}^{2}\sigma_{l}^{2}}(V_{j}-V_{l})^{2}=0,$
since $V_{j}=V_{l}=\operatorname{tr}(\Gamma)$ for all $j,l$. And the location component at the population level simplifies to:
$F=V_{p}-\sum_{l=1}^{g}\lambda_{l}V_{l}=\sum_{l=1}^{g}\lambda_{l}d^{2}(C_{l}^{*},C^{*}),$
representing the weighted sum of squared Riemannian distances from population group means to the overall population mean.
Under contiguous alternatives with $C_{l}^{*}=\exp_{C^{*}}(n^{-1/2}H_{l})$ for symmetric matrices $H_{l}\in T_{C^{*}}\mathrm{SPD}(p)$, we have
$d^{2}(C_{l}^{*},C^{*})=\|n^{-1/2}H_{l}\|_{F}^{2}=n^{-1}\|H_{l}\|_{F}^{2}.$
Thus, the location component becomes
$F=n^{-1}\sum_{l=1}^{g}\lambda_{l}\|H_{l}\|_{F}^{2}.$
The asymptotic distribution of the test statistic $T_{n}=nF_{n}^{2}/\hat{\sigma}^{2}$ then follows from the chi-squared approximation, where $\hat{\sigma}^{2}\to\operatorname{tr}(\Gamma)$, yielding
$T_{n}\xrightarrow{d}\chi^{2}_{g-1}\left(\frac{n\left(\sum_{l=1}^{g}\lambda_{l}\|H_{l}\|_{F}^{2}\right)^{2}}{\operatorname{tr}(\Gamma)\sum_{l=1}^{g}\lambda_{l}^{2}}\right),$
which gives the stated non-centrality parameter. ∎
Proposition 6 (Non-centrality under Alternative Hypothesis).
Under the alternative hypothesis $H_{1}$ with population Fréchet means $C_{1}^{*},\ldots,C_{g}^{*}\in\mathrm{SPD}(p)$ where at least two differ, let $C^{*}$ denote the overall population Fréchet mean and define the population tangent vectors
$v_{l}^{*}=\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(C_{l}^{*}))$
for $l=1,\ldots,g$. Then, as $n\to\infty$ with $n_{l}/n\to\pi_{l}>0$:
  1. (1)
    The test statistic satisfies
    $2\log\Lambda^{*}\xrightarrow{d}\chi^{2}_{d(g-1)}(\delta)$
    where $d=p(p+1)/2$ and the non-centrality parameter is
    $\delta=\lim_{n\to\infty}n\cdot\mathrm{tr}(W^{*-1}B^{*})=\lim_{n\to\infty}n\sum_{l=1}^{g}\pi_{l}(v_{l}^{*})^{T}(W^{*})^{-1}v_{l}^{*}$
    with $B^{*}=\sum_{l=1}^{g}\pi_{l}v_{l}^{*}(v_{l}^{*})^{T}$ and $W^{*}$ the population within-group covariance matrix in the tangent space. This represents the sum of squared Mahalanobis distances of group means from the overall mean in the tangent space.
  2. (2)
    The power function at significance level $\alpha$ is
    $\beta_{n}=P\left(-2\log\Lambda^{*}>\chi^{2}_{d(g-1),1-\alpha}\mid H_{1}\right)\to 1$
    as $n\to\infty$, establishing consistency of the test.
Proof.
Let us analyze the between-group scatter under $H_{1}$. The between-group scatter matrix is
$B=\sum_{l=1}^{g}n_{l}\hat{v}_{l}\hat{v}_{l}^{T}$
Since $\hat{v}_{l}\xrightarrow{p}v_{l}^{*}\neq 0$ for at least some $l$ under $H_{1}$, we have
$\frac{B}{n}=\sum_{l=1}^{g}\frac{n_{l}}{n}\hat{v}_{l}\hat{v}_{l}^{T}\xrightarrow{p}\sum_{l=1}^{g}\pi_{l}v_{l}^{*}(v_{l}^{*})^{T}=B^{*}$
where $B^{*}$ is positive semi-definite with rank $\geq 1$. Thus $B=O_{p}(n)$.
For the within-group scatter, we have the matrix
$W=\sum_{l=1}^{g}\sum_{i=1}^{n_{l}}(v_{li}-\hat{v}_{l})(v_{li}-\hat{v}_{l})^{T}$
where $v_{li}=\operatorname{Vec}_{\hat{C}}(\log_{\hat{C}}(C_{li}))$ is the tangent vector representation of the $i$-th observation from group $l$ at the overall sample mean $\hat{C}$. This converges as $W/n\xrightarrow{p}W^{*}$ where
$W^{*}=\sum_{l=1}^{g}\pi_{l}C_{l}$
with $C_{l}=E[(v_{li}-v_{l}^{*})(v_{li}-v_{l}^{*})^{T}]$ being the population covariance within group $l$. Thus $W=O_{p}(n)$.
We now examine the limiting behavior of Wilks’ Lambda. Since both $W$ and $B$ are $O_{p}(n)$, we can write
$\Lambda^{*}=\frac{|W|}{|W+B|}=\frac{|\tilde{W}|}{|\tilde{W}+\tilde{B}|}$
where $\tilde{W}=W/n$ and $\tilde{B}=B/n$. As $n\to\infty$:
$\Lambda^{*}\xrightarrow{p}\frac{|W^{*}|}{|W^{*}+B^{*}|}=\lambda^{*}<1$
Since $B^{*}>0$ under $H_{1}$, we have $\lambda^{*}<1$ strictly.
This leads us to the asymptotic distribution. The log-likelihood ratio statistic
[…]
where the factor of $n$ arises because we express the determinants in terms of $\tilde{W}=W/n$ and $\tilde{B}=B/n$, which converge to finite limits. Using the expansion $\log|I+A|\approx\mathrm{tr}(A)-\frac{1}{2}\mathrm{tr}(A^{2})+O(\|A\|^{3})$ for the matrix $A=\tilde{W}^{-1}\tilde{B}$, we obtain the leading term $n\cdot\mathrm{tr}(W^{*-1}B^{*})$.
Finally, we establish power and consistency. Since $\delta_{n}=O(n)$, the non-centrality parameter grows linearly with sample size. Therefore, for any fixed critical value $c_{\alpha}$:
$\beta_{n}=P(-2\log\Lambda^{*}>c_{\alpha}\mid H_{1})\to 1$
as $n\to\infty$. ∎
Proposition 7 (Comparison of Non-centrality Parameters).
Consider the Riemannian ANOVA and Fréchet ANOVA tests under the Riemannian normal model with isotropic covariance $\Gamma=\sigma^{2}I_{d^{2}}$ and balanced design ($n_{l}=n/g$ for all groups). For a specific alternative where only group 1 deviates from the common mean by Riemannian distance $\theta>0$ (i.e., $d(C_{1}^{*},C^{*})=\theta$ and $d(C_{l}^{*},C^{*})=0$ for $l=2,\ldots,g$), the non-centrality parameters simplify to:
(8) $\delta_{\text{Novel}}$ $=\frac{n\theta^{2}}{g\sigma^{2}}=O(\theta^{2})$
(9) $\delta_{\text{Fréchet}}$ $=\frac{n\theta^{4}}{g\sigma^{2}d^{2}}=O(\theta^{4})$
Proof.
We establish the non-centrality expressions for both tests under the specified conditions.
The general expression for the non-centrality parameter can be rewritten as:
$\delta_{\text{Novel}}=n\sum_{l=1}^{g}\lambda_{l}\|v_{l}^{*}\|_{\Gamma^{-1}}^{2}$
where $v_{l}^{*}=\operatorname{Vec}(\log_{C^{*}}(C_{l}^{*}))$ and $\lambda_{l}=1/g$ for balanced designs.
For the specific alternative with only group 1 deviating:
$\|v_{1}^{*}\|_{\Gamma^{-1}}^{2}=\frac{1}{\sigma^{2}}\|v_{1}^{*}\|_{F}^{2}=\frac{\theta^{2}}{\sigma^{2}}$
Therefore:
$\delta_{\text{Novel}}=\frac{n\theta^{2}}{g\sigma^{2}}$
On the other hand, the Fréchet ANOVA non-centrality parameter is:
$\delta_{\text{Fréchet}}=\frac{n\left(\sum_{l=1}^{g}\lambda_{l}d^{2}(C_{l}^{*},C^{*})\right)^{2}}{\operatorname{tr}(\Gamma)\sum_{l=1}^{g}\lambda_{l}^{2}}$
Substituting these values:
$\delta_{\text{Fréchet}}=\frac{n\cdot(\theta^{2}/g)^{2}}{\sigma^{2}d^{2}\cdot(1/g)}=\frac{n\theta^{4}}{g\sigma^{2}d^{2}}$
Remark 1 (Practical Implications).
The quadratic versus quartic dependence on $\theta$ has significant practical consequences. For detecting small location shifts, the Riemannian ANOVA test requires substantially smaller sample sizes to achieve the same power as Fréchet ANOVA. However, Fréchet ANOVA retains advantages in robustness to model misspecification and the ability to detect variance heterogeneity through its scale component $U_{n}$. The choice between tests should consider both the expected effect size and confidence in the Riemannian normality assumption.
3. Conclusion
We have developed a novel framework for statistical inference on brain connectivity data that leverages the full geometric structure of the symmetric positive definite manifold. Our approach addresses fundamental limitations of existing methods by introducing a Riemannian generalization of classical MANOVA that properly exploits the dimensionality of connectome data.
In the methods section, we introduced three key contributions. First, we developed the Mahalanobis Affine Invariant Riemannian Metric (MAIRM), which relaxes the sphericity assumptions inherent in existing Riemannian models while maintaining computational tractability. This framework provides consistent estimators for both location and dispersion parameters and reduces to classical multivariate analysis when applied to Euclidean data. Second, we constructed a test statistic that generalizes Wilks’ Lambda to the SPD manifold, establishing a rigorous connection between Euclidean and Riemannian inference. Our statistic respects the manifold geometry while preserving the interpretability of classical MANOVA. Third, we derived the asymptotic distribution of our test statistic under both null and alternative hypotheses, enabling practical hypothesis testing with theoretical guarantees.
The theoretical analysis revealed a striking advantage of our approach over existing methods. While Fréchet ANOVA’s non-centrality parameter scales as $O(\theta^{4})$ in the effect size, our Riemannian ANOVA achieves $O(\theta^{2})$ scaling. This quadratic versus quartic dependence translates to substantially improved power for detecting location differences, particularly for small to moderate effect sizes commonly encountered in neuroimaging studies. However, this power gain comes at the cost of stronger distributional assumptions—our method requires the data to follow a Riemannian normal distribution, whereas Fréchet ANOVA applies to general metric spaces.
Several directions for future work emerge from this foundation. First, developing goodness-of-fit tests for the Riemannian normal model would enable practitioners to assess whether our stronger assumptions are justified for their data. Second, extending our framework to handle heterogeneous covariance structures across groups would broaden its applicability. Third, incorporating covariates and regression structures while maintaining the geometric insights developed here represents an important direction for more flexible modeling. Finally, computational optimizations exploiting the exact formulas derived in our appendices could enable application to high-dimensional connectome data from modern neuroimaging studies.
The framework presented here demonstrates that carefully exploiting manifold structure can yield substantial gains in statistical efficiency. By moving beyond distance-based methods to embrace the full geometric toolkit available on Riemannian manifolds, we open new possibilities for powerful and interpretable inference on complex structured data.
Riemannian Geometry Details
For a parcellation of size $p$, connectomes are symmetric, positive definite, $p\times p$ matrices. The set $\mathcal{M}=\mathrm{SPD}_{p}$ is a convex cone, but it is not a linear subspace of the set of $p\times p$ matrices, $\mathrm{Mat}\left(p\right)$. They are an open subset of the linear subspace of $p\times p$ symmetric matrices, $\mathrm{Symm}_{p}$. As such, they constitute a smooth manifold of dimension $d=\frac{p\left(p+1\right)}{2}$. We denote a generic point of $\mathcal{M}$ as $C$. The tangent space at any $C$, $T_{C}\mathcal{M}$, is naturally isomorphic to $\mathrm{Symm}_{p}$.
Furthermore, $\mathcal{M}$ can be endowed with the structure of a Riemannian manifold by choosing a Riemannian metric. As mentioned before, a popular choice is AIRM (pennec). It can be introduced in a couple of ways. On one hand, it is a Fisher information metric related to the fact that each $C\in\mathcal{M}$ is the set of parameters for a distribution, namely $N\left(0,C\right)$ (lenglet). On the other hand, $\mathcal{M}$ is a homogeneous space of the Lie group $\mathrm{Mat}\left(p\right)^{\times}$ under the action given by $C\mapsto ACA^{T}$ for $A\in\mathrm{Mat}\left(p\right)^{\times}$. The left invariant metric on $\mathrm{Mat}\left(p\right)^{\times}$ descends to a metric on $\mathcal{M}$.
AIRM has a prominent role in part for historic reasons. But it has other nice properties. It is a complete metric: geodesics stay on the manifold for infinite time (the boundary of $\mathcal{M}$, the set of semidefinite matrices, are at infinite distance in AIRM). Also, it has negative curvature. These two facts allow us to use Hadamard’s theorem to conclude that for each $C$, the maps $\exp_{C}:\mathcal{M}\rightarrow T_{C}\mathcal{M}$ and $\log_{C}:T_{C}\mathcal{M}\rightarrow\mathcal{M}$ are global, mutually inverse diffeomorphisms. AIRM introduces a geodesic distance $d\left(\cdot,\cdot\right)$ on $\mathcal{M}$ as well as a norm $\norm{\cdot}_{C}$ on $T_{C}\mathcal{M}$, which satisfy the relation
$\norm{\log_{C}\left(\Lambda\right)}_{C}=d\left(C,\Lambda\right)$
AIRM allows us to introduce several probabilistic features which we can use to perform statistical inference on them. First, it allows us to compute expected values. We cannot perform integration due to the lack of a linear structure. Instead, we use the notion of Fréchet mean: for a random connectome $C$,
$\mathbb{E}\left[C\right]=\operatorname{argmin}_{\Lambda}\norm{\log_{\Lambda}\left(C\right)}_{\Lambda}^{2}$
We can also introduce a notion of normal distribution, as follows. AIRM provides us with a linear isometry $\operatorname{Vec}_{C}:T_{C}\mathcal{M}\rightarrow\mathbb{R}^{p}$. For matrices $C^{\ast}\in\mathcal{M},\Gamma\in\mathrm{SPD}_{d}$, we say the random connectome $C$ has distribution $N\left(C^{\ast},\Gamma\right)$ if the random vector $\operatorname{Vec}_{C^{\ast}}\left(\log_{C^{\ast}}\left(C\right)\right)$ has distribution $N\left(0,\Gamma\right)$.
Theorem 8 (Jacobi Field Quotient Identity).
Let $M$ be a Riemannian manifold and $\gamma:[0,1]\to M$ be the geodesic $\gamma(s)=\exp_{C}(s\log_{C}(A))$ connecting $C$ to $A$. Consider a variation $\sigma(s,t)$ with:
  • $\sigma(0,t)=C$ for all $t$
  • $\sigma(s,0)=\gamma(s)$
  • $\sigma(1,t)=B(t)$ where $B(0)=B$
Let $J(s)=\frac{\partial\sigma}{\partial t}\big|_{t=0}$ be the associated Jacobi field. Then for small $\varepsilon>0$:
$\frac{\log_{\gamma(\varepsilon)}(B)-\Pi_{C\to\gamma(\varepsilon)}\left[\log_{C}(B)\right]}{\varepsilon}=\nabla_{s}J\bigg|_{s=\varepsilon}+O(\varepsilon)$
where $\Pi_{C\to\gamma(\varepsilon)}$ denotes parallel transport along $\gamma$ from $C$ to $\gamma(\varepsilon)$.
Proof.
We establish the result through a careful analysis of the Jacobi field evolution and its covariant derivatives.
Let $\{E_{i}(s)\}_{i=1}^{n}$ be a parallel orthonormal frame along $\gamma$, constructed by parallel transporting an orthonormal basis from $T_{C}M$ along the geodesic. Since the frame is parallel, we have $\nabla_{s}E_{i}=0$ for all $i$ and $s\in[0,1]$.
Express the Jacobi field in this frame:
$J(s)=\sum_{i=1}^{n}J^{i}(s)E_{i}(s)$
where $J^{i}(s)=\langle J(s),E_{i}(s)\rangle$ are the component functions.
Since the frame is parallel, the covariant derivatives of $J$ have a simple form:
$\nabla_{s}J=\sum_{i=1}^{n}\frac{dJ^{i}}{ds}E_{i}(s),\quad\nabla_{s}^{2}J=\sum_{i=1}^{n}\frac{d^{2}J^{i}}{ds^{2}}E_{i}(s)$
The Jacobi equation $\nabla_{s}^{2}J+R(J,\dot{\gamma})\dot{\gamma}=0$ becomes:
$\sum_{i=1}^{n}\frac{d^{2}J^{i}}{ds^{2}}E_{i}(s)+R(J(s),\dot{\gamma}(s))\dot{\gamma}(s)=0$
In the parallel frame, the curvature endomorphism $R(\cdot,\dot{\gamma})\dot{\gamma}$ has components:
$R(J,\dot{\gamma})\dot{\gamma}=\sum_{i,j=1}^{n}R_{ij}J^{j}E_{i}$
where $R_{ij}=\langle R(E_{j},\dot{\gamma})\dot{\gamma},E_{i}\rangle$.
Therefore, the component equations are:
$\frac{d^{2}J^{i}}{ds^{2}}+\sum_{j=1}^{n}R_{ij}(s)J^{j}(s)=0,\quad i=1,\ldots,n$
For each component function $J^{i}(s)$, we have the Taylor expansion around $s=0$:
$J^{i}(s)=J^{i}(0)+s\frac{dJ^{i}}{ds}\bigg|_{s=0}+\frac{s^{2}}{2}\frac{d^{2}J^{i}}{ds^{2}}\bigg|_{s=0}+O(s^{3})$
From the Jacobi equation at $s=0$:
$\frac{d^{2}J^{i}}{ds^{2}}\bigg|_{s=0}=-\sum_{j=1}^{n}R_{ij}(0)J^{j}(0)$
Substituting:
$J^{i}(\varepsilon)=J^{i}(0)+\varepsilon\frac{dJ^{i}}{ds}\bigg|_{s=0}-\frac{\varepsilon^{2}}{2}\sum_{j=1}^{n}R_{ij}(0)J^{j}(0)+O(\varepsilon^{3})$
The parallel transport of $J(0)\in T_{C}M$ to $T_{\gamma(\varepsilon)}M$ is given by:
$\Pi_{C\to\gamma(\varepsilon)}[J(0)]=\sum_{i=1}^{n}J^{i}(0)E_{i}(\varepsilon)$
This vector in $T_{\gamma(\varepsilon)}M$ has components $J^{i}(0)$ in the parallel frame at $\gamma(\varepsilon)$.
The Jacobi field at $\gamma(\varepsilon)$ is:
$J(\varepsilon)=\sum_{i=1}^{n}J^{i}(\varepsilon)E_{i}(\varepsilon)$
Both $J(\varepsilon)$ and $\Pi_{C\to\gamma(\varepsilon)}[J(0)]$ are vectors in $T_{\gamma(\varepsilon)}M$. Their difference is:
$J(\varepsilon)-\Pi_{C\to\gamma(\varepsilon)}[J(0)]=\sum_{i=1}^{n}[J^{i}(\varepsilon)-J^{i}(0)]E_{i}(\varepsilon)$
Substituting the expansion:
$J(\varepsilon)-\Pi_{C\to\gamma(\varepsilon)}[J(0)]=\varepsilon\sum_{i=1}^{n}\frac{dJ^{i}}{ds}\bigg|_{s=0}E_{i}(\varepsilon)-\frac{\varepsilon^{2}}{2}\sum_{i,j=1}^{n}R_{ij}(0)J^{j}(0)E_{i}(\varepsilon)+O(\varepsilon^{3})$
Note that $\sum_{i=1}^{n}\frac{dJ^{i}}{ds}\big|_{s=0}E_{i}(\varepsilon)$ is the parallel transport of $\nabla_{s}J\big|_{s=0}$ to $T_{\gamma(\varepsilon)}M$.
Dividing by $\varepsilon$:
$\frac{J(\varepsilon)-\Pi_{C\to\gamma(\varepsilon)}[J(0)]}{\varepsilon}=\sum_{i=1}^{n}\frac{dJ^{i}}{ds}\bigg|_{s=0}E_{i}(\varepsilon)-\frac{\varepsilon}{2}\sum_{i,j=1}^{n}R_{ij}(0)J^{j}(0)E_{i}(\varepsilon)+O(\varepsilon^{2})$
The covariant derivative $\nabla_{s}J$ along the geodesic has components $\nabla_{s}J]^{i}(s)=\frac{dJ^{i}}{ds}(s)$.
From the Jacobi equation:
$\frac{d}{ds}\left[\frac{dJ^{i}}{ds}\right]=\frac{d^{2}J^{i}}{ds^{2}}=-\sum_{j=1}^{n}R_{ij}(s)J^{j}(s)$
For small $\varepsilon$:
$\nabla_{s}J]^{i}(\varepsilon)=[\nabla_{s}J]^{i}(0)-\varepsilon\sum_{j=1}^{n}R_{ij}(0)J^{j}(0)+O(\varepsilon^{2})$
Therefore, $\nabla_{s}J\big|_{s=\varepsilon}$ expressed in the frame at $\gamma(\varepsilon)$ is:
$\nabla_{s}J\big|_{s=\varepsilon}=\Pi_{C\to\gamma(\varepsilon)}[\nabla_{s}J\big|_{s=0}]-\varepsilon\Pi_{C\to\gamma(\varepsilon)}[R(J(0),\dot{\gamma}(0))\dot{\gamma}(0)]+O(\varepsilon^{2})$
From the quotient expression and the evolution of $\nabla_{s}J$:
$\Pi_{C\to\gamma(\varepsilon)}[\nabla_{s}J\big|_{s=0}]=\nabla_{s}J\big|_{s=\varepsilon}+\varepsilon\Pi_{C\to\gamma(\varepsilon)}[R(J(0),\dot{\gamma}(0))\dot{\gamma}(0)]+O(\varepsilon^{2})$
Substituting:
$\frac{J(\varepsilon)-\Pi_{C\to\gamma(\varepsilon)}[J(0)]}{\varepsilon}=\nabla_{s}J\big|_{s=\varepsilon}+\frac{\varepsilon}{2}\Pi_{C\to\gamma(\varepsilon)}[R(J(0),\dot{\gamma}(0))\dot{\gamma}(0)]+O(\varepsilon^{2})$
The curvature term $\frac{\varepsilon}{2}\Pi_{C\to\gamma(\varepsilon)}[R(J(0),\dot{\gamma}(0))\dot{\gamma}(0)]$ is of order $\varepsilon$, so:
$\frac{J(\varepsilon)-\Pi_{C\to\gamma(\varepsilon)}[J(0)]}{\varepsilon}=\nabla_{s}J\big|_{s=\varepsilon}+O(\varepsilon)$
Finally, since $\gamma(\varepsilon)=\exp_{C}(\varepsilon\log_{C}(A))$ and $J(\varepsilon)=\log_{\gamma(\varepsilon)}(B)$ and $J(0)=\log_{C}(B)$:
$\frac{\log_{\gamma(\varepsilon)}(B)-\Pi_{C\to\gamma(\varepsilon)}[\log_{C}(B)]}{\varepsilon}=\nabla_{s}J\bigg|_{s=\varepsilon}+O(\varepsilon)$
Corollary 9 (Logarithm Difference Formula).
Let $C,A,B\in M$ and consider the geodesic $\gamma(s)=\exp_{C}(s\log_{C}(A))$. For small $\varepsilon>0$, let $\gamma(\varepsilon)$ be a point on this geodesic close to $C$. Then:
$\frac{\log_{\gamma(\varepsilon)}(B)-\Pi_{C\to\gamma(\varepsilon)}[\log_{C}(B)-\log_{C}(A)]-\log_{\gamma(\varepsilon)}(A)}{\varepsilon}=\nabla_{s}\tilde{J}\bigg|_{s=\varepsilon}+O(\varepsilon)$
where $\tilde{J}(s)$ is the Jacobi field along $\gamma$ with initial conditions:
  • $\tilde{J}(0)=\log_{C}(B)-\log_{C}(A)$
  • $\tilde{J}(1)=0$ (since $\sigma(1,t)=A$ for all $t$ in the variation)
Proof.
Consider the variation $\sigma(s,t)$ defined by:
$\sigma(s,t)=\exp_{\gamma(s)}(t[\log_{\gamma(s)}(B)-\log_{\gamma(s)}(A)])$
This variation satisfies: $\sigma(s,0)=\gamma(s)$ for all $s$, $\sigma(0,t)=\exp_{C}(t[\log_{C}(B)-\log_{C}(A)])$, and $\sigma(1,t)=A$ for all $t$ (since $\gamma(1)=A$ and $\log_{A}(A)=0$).
The associated Jacobi field is:
$\tilde{J}(s)=\frac{\partial\sigma}{\partial t}\bigg|_{t=0}=\log_{\gamma(s)}(B)-\log_{\gamma(s)}(A)$
Now we can decompose:
$\log_{\gamma(\varepsilon)}(B)=[\log_{\gamma(\varepsilon)}(B)-\log_{\gamma(\varepsilon)}(A)]+\log_{\gamma(\varepsilon)}(A)$
The first term is $\tilde{J}(\varepsilon)$, which by Theorem 8 equals:
$\tilde{J}(\varepsilon)=\Pi_{C\to\gamma(\varepsilon)}[\tilde{J}(0)]+\varepsilon\nabla_{s}\tilde{J}\big|_{s=\varepsilon}+O(\varepsilon^{2})$
Since $\tilde{J}(0)=\log_{C}(B)-\log_{C}(A)$:
$\log_{\gamma(\varepsilon)}(B)-\log_{\gamma(\varepsilon)}(A)=\Pi_{C\to\gamma(\varepsilon)}[\log_{C}(B)-\log_{C}(A)]+\varepsilon\nabla_{s}\tilde{J}\big|_{s=\varepsilon}+O(\varepsilon^{2})$
Rearranging and dividing by $\varepsilon$:
$\frac{\log_{\gamma(\varepsilon)}(B)-\Pi_{C\to\gamma(\varepsilon)}[\log_{C}(B)-\log_{C}(A)]-\log_{\gamma(\varepsilon)}(A)}{\varepsilon}=\nabla_{s}\tilde{J}\bigg|_{s=\varepsilon}+O(\varepsilon)$
Lemma 10 (Derivative of Logarithm Map with Moving Base Point).
For a curve $\gamma(s)$ in the manifold of SPD matrices with velocity $\dot{\gamma}(s)$ and a fixed point $B$, the derivative of the logarithm map is:
$\frac{d}{ds}\bigg|_{s=0}\log_{\gamma(s)}(B)=-\nabla_{\dot{\gamma}(0)}\log_{\gamma(0)}(B)$
where $\nabla$ is the Levi-Civita connection of the AIRM metric.
Proof.
This follows from the differential geometry of the logarithm map. The logarithm map $\log:\mathcal{M}\times\mathcal{M}\to T\mathcal{M}$ can be viewed as a map $(P,Q)\mapsto\log_{P}(Q)\in T_{P}\mathcal{M}$. When we vary the base point $P$ along a curve $\gamma(s)$ while keeping $Q=B$ fixed, the change in $\log_{\gamma(s)}(B)$ involves the intrinsic change due to moving the base point (captured by the connection) and the parallel transport of the vector as we move along $\gamma$. The formula $\nabla_{\dot{\gamma}(0)}\log_{\gamma(0)}(B)$ encodes precisely this geometric change. ∎
Proposition 11 (AIRM Jacobi Field Velocity).
Let $(\mathcal{M},g^{\mathrm{AIRM}})$ be the manifold of $n\times n$ symmetric positive definite matrices equipped with the Affine Invariant Riemannian Metric (AIRM). Let $C,A,B\in\mathcal{M}$ be three points, and let $\gamma:[0,1]\to\mathcal{M}$ be the geodesic from $C$ to $A$, given by:
$\gamma(s)=C^{1/2}(C^{-1/2}AC^{-1/2})^{s}C^{1/2}$
Define the Jacobi field $J(s)$ along $\gamma$ arising from the variation:
$J(s)=\log_{\gamma(s)}(B)-\log_{\gamma(s)}(A)$
Then the velocity of the Jacobi field at $s=0$ is given exactly by:
$\dot{J}(0)=-\frac{1}{2}[\log_{C}(A),\log_{C}(B)]-\frac{1}{2}\{\log_{C}(A),\log_{C}(B)\}_{C}+\log_{C}(A)$
where:
  • [X,Y]=XYYX[X,Y]=XY-YX is the matrix commutator
  • $\{X,Y\}_{C}=C^{-1}XY+YX^{T}C^{-1}$ is the symmetrized product with the metric at $C$
  • $\log_{P}(Q)=P^{1/2}\log(P^{-1/2}QP^{-1/2})P^{1/2}$ is the Riemannian logarithm map
Proof.
We denote $L_{A}=\log(C^{-1/2}AC^{-1/2})$ as the matrix logarithm, $L_{B}=\log(C^{-1/2}BC^{-1/2})$, and $M=C^{-1/2}AC^{-1/2}$ as the conjugated target matrix. Then $\log_{C}(A)=C^{1/2}L_{A}C^{1/2}$, $\log_{C}(B)=C^{1/2}L_{B}C^{1/2}$, $\gamma(s)=C^{1/2}M^{s}C^{1/2}$, and $\dot{\gamma}(s)=C^{1/2}L_{A}M^{s}C^{1/2}$.
First, we compute $\frac{d}{ds}\log_{\gamma(s)}(A)$. Since $\gamma(s)$ is the geodesic from $C$ to $A$ with $\gamma(0)=C$ and $\gamma(1)=A$, the remaining distance along the geodesic from $\gamma(s)$ to $A$ is (1s)(1-s) times the initial distance. By the properties of geodesics in Riemannian manifolds:
$\log_{\gamma(s)}(A)=(1-s)\log_{C}(A)$
Therefore:
$\frac{d}{ds}\log_{\gamma(s)}(A)=-\log_{C}(A)$
Next, we compute $\frac{d}{ds}\bigg|_{s=0}\log_{\gamma(s)}(B)$ using Lemma 10. For the AIRM metric at point $C\in\mathcal{M}$, the Levi-Civita connection is:
$\nabla_{X}Y=DY[X]-\frac{1}{2}(C^{-1}XY+YX^{T}C^{-1})$
where $X,Y\in T_{C}\mathcal{M}$ are tangent vectors (symmetric matrices) and $DY[X]$ is the directional derivative.
Applying the lemma with $\dot{\gamma}(0)=\log_{C}(A)$:
$\frac{d}{ds}\bigg|_{s=0}\log_{\gamma(s)}(B)=-\nabla_{\log_{C}(A)}\log_{C}(B)$
Since $\log_{C}(A)$ and $\log_{C}(B)$ are constant vector fields when viewed in the tangent space at $C$, the directional derivative term vanishes: $D(\log_{C}(B))[\log_{C}(A)]=0$.
The covariant derivative becomes:
$\nabla_{\log_{C}(A)}\log_{C}(B)=-\frac{1}{2}(C^{-1}\log_{C}(A)\log_{C}(B)+\log_{C}(B)\log_{C}(A)^{T}C^{-1})$
Since $\log_{C}(A)$ and $\log_{C}(B)$ are symmetric matrices, we have $\log_{C}(A)^{T}=\log_{C}(A)$. We can decompose the product of two matrices into symmetric and anti-symmetric parts:
$\log_{C}(A)\log_{C}(B)=\frac{1}{2}[\log_{C}(A),\log_{C}(B)]+\frac{1}{2}\{\log_{C}(A),\log_{C}(B)\}$
where $\{X,Y\}=XY+YX$ is the anticommutator. After careful computation:
$\nabla_{\log_{C}(A)}\log_{C}(B)=\frac{1}{2}[\log_{C}(A),\log_{C}(B)]+\frac{1}{2}\{\log_{C}(A),\log_{C}(B)\}_{C}$
Therefore:
$\frac{d}{ds}\bigg|_{s=0}\log_{\gamma(s)}(B)=-\frac{1}{2}[\log_{C}(A),\log_{C}(B)]-\frac{1}{2}\{\log_{C}(A),\log_{C}(B)\}_{C}$
Finally, the velocity of the Jacobi field is:
$\dot{J}(0)=\frac{d}{ds}\bigg|_{s=0}[\log_{\gamma(s)}(B)-\log_{\gamma(s)}(A)]$
$\left[-\frac{1}{2}[\log_{C}(A),\log_{C}(B)]-\frac{1}{2}\{\log_{C}(A),\log_{C}(B)\}_{C}\right]-(-\log_{C}(A))$
Thus:
$\dot{J}(0)=-\frac{1}{2}[\log_{C}(A),\log_{C}(B)]-\frac{1}{2}\{\log_{C}(A),\log_{C}(B)\}_{C}+\log_{C}(A)$
Corollary 12 (Covariant Derivative of Jacobi Field).
Under the same conditions as Proposition 11, the covariant derivative of the Jacobi field at $s=0$ is:
$\nabla_{s}J|_{s=0}=\dot{J}(0)-\frac{1}{2}(C^{-1}\dot{\gamma}(0)J(0)+J(0)\dot{\gamma}(0)^{T}C^{-1})$
where $J(0)=\log_{C}(B)-\log_{C}(A)$ and $\dot{\gamma}(0)=\log_{C}(A)$.
Remark 2.
The formula in Proposition 11 is exact, not an approximation. All terms can be computed using standard matrix operations: matrix logarithms (computable via eigendecomposition), matrix multiplications, and additions. No differential equation solving is required, making the computation both efficient and numerically stable.
The geometric interpretation is clear: the term $\log_{C}(A)$ represents the linear interpolation ensuring $J(1)=0$, the commutator $\log_{C}(A),\log_{C}(B)]$ captures the curvature effect from non-commutativity, and the symmetric term $\{\log_{C}(A),\log_{C}(B)\}_{C}$ arises from the metric structure through the Christoffel symbols.
Tables, Listings and Figures