### v2

# A Novel Testing Approach for Differences Among Brain Connectomes

Abstract
The study of samples that take values in non-Euclidean metric spaces is an important part of modern statistics literature. For instance, it includes brain connectomics, an active area of neuroscience where data consists of symmetric, positive definite (SPD) matrices. Now, Fréchet ANOVA has played a significant role in the study of non-Euclidean samples, due in part to its wide applicability. This paper addresses two gaps in that literature that derive from two observations about that method. First, that it is not a generalization of MANOVA; in fact, such a generalization has not been sufficiently explored, and our first goal is to propose one. Second, that its power function has not been systematically studied; our second goal is to shed some light on that subject, including an argument for how in some scenarios our generalization is more powerful. Hopefully, these contributions will demonstrate that there are additional directions in which the study of non-Euclidean samples can go and hence spark more interest in the subject.
Riemannian geometry,
keywords:
[…]
1 Introduction
The goal of generalizing statistical methods, particularly multivariate ones, to settings in which data poses structure other than (or on top of) a regular vector is an old one. The practice of statistics requires the measure of similarity, so these generalizations are often made in the setting of metric spaces. Underlying probabilistic methods for that setting have been based on the concept of Fréchet mean (Karcher, 1977). The sample mean has been shown to possess desirable properties including existence, uniqueness for negatively curved manifolds (Kendall, 1990), consistency (Bhattacharya and Patrangenaru, 2003), and asymptotic normality (Hotz, Le and Wood, 2024). Using that concept, we can generalize traditional statistical models like linear regression to that setting as well (Thomas Fletcher, 2013).
Oftentimes, data has a layer of structure that sits on top of the metric space, that of a Riemannian manifold, allowing for additional complexity in statistical modeling, strengthening what we can say about linear regression (Cornea et al., 2017) and facilitating the construction of other models.
We obtain yet another layer of structure when we focus on specific manifolds, the one of symmetric positive definite $p\times p$ matrices, $\mathcal{M}=\mathrm{SPD}_{p}$ being the one that is of interest to us. The seminal work of several researchers (Pennec, Fillard and Ayache, 2006; Arsigny et al., 2005; Lenglet et al., 2006a) illustrate how the choice of a specific Riemannian metric turns these abstract ideas into actually computable statistical models with concrete applications to brain imaging (Varoquaux et al., 2010). This hierarchy of structure means that statistical modeling of non-Euclidean data lives on a spectrum. On the general side of the spectrum, if we make mild assumptions about the structure of the data, we obtain widely applicable models. On the specific side, if we have a concrete structure in mind, we can use additional tools.
Fréchet ANOVA (Dubey and Müller, 2019) is a seminal statistical test for comparing multiple samples of non-Euclidean data. It sits on the general side of that spectrum; it only assumes a metric space structure with mild conditions. The goal of this paper is to propose and analyze a test for a similar hypothesis on the specific side of the spectrum. The idea behind this new test is that the properties of certain Riemannian metrics on $\mathcal{M}$ are close enough to those of Euclidean space as to allow for classical ANOVA to carry over. In this way, we obtain a test that is much less general than Fréchet ANOVA, but as we show, what we lose in generality we gain in power, at least in a certain region of the parameter space.
On a more technical level, this paper addresses two big challenges. First, carrying over classical ANOVA requires us to identify a way to obtain the decomposition of squares that characterizes that method in the manifold setting. We do so by studying projections of the sample to the tangent space of the Fréchet mean. Second, we identify distributional assumptions under which we can derive an asymptotic distribution of the new test statistic. Those assumptions correspond to what we call the Riemannian normal likelihood. We also observe that said likelihood has a Riemannian metric associated to it, which we call the MAIRM.
The remainder of this paper is organized as follows. In Section 2, we develop our methodological framework in three stages. We begin by introducing parameter estimation for the MAIRM, a generalization of classical affine invariant metric that relaxes sphericity assumptions. We then construct our Riemannian generalization of Wilks’ Lambda, establishing its fundamental properties and showing how it reduces to classical multivariate ANOVA in the Euclidean case. Finally, we derive the asymptotic distribution of our test statistic under the null hypothesis and analyze its asymptotic behavior under alternatives, formally comparing its power properties to those of Fréchet ANOVA. Section 3 summarizes our contributions. Technical details on the Riemannian geometry of the SPD manifold and additional computational formulas are provided in the appendix.
2 Contributions
2.1 Mahalanobis AIRM
In a number of statistical models based on Riemannian geometry ideas, it is assumed that the likelihood depends exclusively on the geodesic distance between observations and some reference point. The simplest such model is decribed in Varoquaux et al. (2010): for a manifold-valued random variable $\boldsymbol{C}$,
$p(\boldsymbol{C})\propto\exp\left(-\frac{1}{2\sigma^{2}}\left\lVert\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C})\right\rVert_{\boldsymbol{C}^{\ast}}^{2}\right)$
for some $\boldsymbol{C}^{\ast}\in\mathcal{M},\sigma\in\mathbb{R}^{+}$. Here, $\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C})$ denotes the Riemannian logarithm map. Inference for such a model is simple: given an iid sample $\{\boldsymbol{C}_{i}\}_{i}$, the MLE estimators for $\boldsymbol{C}^{\ast}$ and $\sigma^{2}$ are the sample Fréchet mean, $\bar{\boldsymbol{C}}_{n}$ and
$\hat{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}\left\lVert\log_{\bar{\boldsymbol{C}}_{n}}(\boldsymbol{C}_{i})\right\rVert_{\bar{\boldsymbol{C}}_{n}}^{2}$
These estimators are independent.
Such a model makes a sphericity assumption. To make it explicit, consider the AIRM metric for concreteness and recall the vectorization maps (Pennec, Fillard and Ayache, 2006):
$\operatorname{Vect}_{I}:T_{I}\mathcal{M}$ $\rightarrow\mathbb{R}^{p(p+1)/2}$
$\{w_{ij}\}$ $\mapsto(w_{11},\sqrt{2}w_{12},w_{22},\ldots)$
and
$\operatorname{Vect}_{\boldsymbol{C}}:T_{\boldsymbol{C}}\mathcal{M}$ $\rightarrow\mathbb{R}^{p(p+1)/2}$
$\boldsymbol{\Lambda}$ $\mapsto\operatorname{Vect}_{I}(\boldsymbol{C}^{-1/2}\boldsymbol{\Lambda}\boldsymbol{C}^{-1/2})$
Here we have used the identification of $T_{\boldsymbol{C}}\mathcal{M}$ with the vector space of $p$ by $p$ symmetric matrices. Both maps are linear isometries.
The model can be rewritten using these maps as
$p(\boldsymbol{C})\propto\exp\left(-\frac{1}{2\sigma^{2}}\operatorname{Vect}_{\boldsymbol{C}^{\ast}}(\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C}))^{T}\operatorname{Vect}_{\boldsymbol{C}^{\ast}}(\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C}))\right)$
In other words $\operatorname{Vect}_{\boldsymbol{C}^{\ast}}(\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C}))\sim N_{p(p+1)/2}(0,\sigma^{2}I)$.
Let us relax that assumption and consider the model
$p(\boldsymbol{C})\propto\exp\left(-\frac{1}{2}\operatorname{Vect}_{\boldsymbol{C}^{\ast}}(\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C}))^{T}\Gamma^{-1}\operatorname{Vect}_{\boldsymbol{C}^{\ast}}(\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C}))\right)$
for some positive definite $\Gamma$.
Given an iid sample $\{\boldsymbol{C}_{i}\}$, it is not hard to come up with natural estimators for $\boldsymbol{C}^{\ast}$ and $\Gamma$. First, notice that $\boldsymbol{C}^{\ast}$ is still the population Fréchet mean, so the sample Fréchet mean $\bar{\boldsymbol{C}}_{n}$ can still be used to estimate it. As for $\Gamma$, the estimator would be
$\hat{\mathbf{S}}_{n}=\frac{1}{n-1}\sum_{i=1}^{n}\operatorname{Vect}_{\bar{\boldsymbol{C}}_{n}}(\log_{\bar{\boldsymbol{C}}_{n}}(\boldsymbol{C}_{i}))\operatorname{Vect}_{\bar{\boldsymbol{C}}_{n}}(\log_{\bar{\boldsymbol{C}}_{n}}(\boldsymbol{C}_{i}))^{T}$
The price we pay for the added generality of this model is that it is no longer clear that $\bar{\boldsymbol{C}}_{n}$ and $(n-1)/n\hat{\mathbf{S}}_{n}$ are MLE estimators for $\boldsymbol{C}^{\ast}$ and $\Gamma$.
However, the fact that $\bar{\boldsymbol{C}}_{n}$ converges in probability to $\boldsymbol{C}^{\ast}$ is a direct consequence of Theorem 2.3 in Bhattacharya and Patrangenaru (2005). In addition:
Proposition 2.1.
$\hat{\mathbf{S}}_{n}$ converges in probability to $\Gamma$.
Proof.
We have the chain of implications:
$\bar{\boldsymbol{C}}_{n}\overset{P}{\rightarrow}\boldsymbol{C}^{\ast}$
$\Rightarrow\bar{\boldsymbol{C}}_{n}^{-1/2}\boldsymbol{C}_{i}\bar{\boldsymbol{C}}_{n}^{-1/2}\overset{P}{\rightarrow}(\boldsymbol{C}^{\ast})^{-1/2}\boldsymbol{C}_{i}(\boldsymbol{C}^{\ast})^{-1/2}$
$\Rightarrow\log(\bar{\boldsymbol{C}}_{n}^{-1/2}\boldsymbol{C}_{i}\bar{\boldsymbol{C}}_{n}^{-1/2})\overset{P}{\rightarrow}$
$\qquad\log((\boldsymbol{C}^{\ast})^{-1/2}\boldsymbol{C}_{i}(\boldsymbol{C}^{\ast})^{-1/2})$
$\Rightarrow\operatorname{Vect}_{\bar{\boldsymbol{C}}_{n}}(\log_{\bar{\boldsymbol{C}}_{n}}(\boldsymbol{C}_{i}))\overset{P}{\rightarrow}\operatorname{Vect}_{\boldsymbol{C}^{\ast}}(\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C}_{i}))$
from which the result follows easily. ∎
A sphericity assumption is also used in regression models with Euclidean predictors and connectome-valued responses. A popular model (Thomas Fletcher, 2013) is based on choosing $\boldsymbol{C}^{\ast}\in\mathcal{M}$, $\{v_{k}\}_{k}\in T_{\boldsymbol{C}^{\ast}}\mathcal{M}$, $\sigma^{2}\in\mathbb{R}^{+}$ and assumming that the observations $\{(\mathbf{x}_{i},\boldsymbol{C}_{i})\}_{i}$ are distributed according to the likelihood
$p(\boldsymbol{C}_{i})\propto\exp\left(-\frac{1}{2\sigma^{2}}\left\lVert\log_{\boldsymbol{C}_{i}^{\ast}}(\boldsymbol{C}_{i})\right\rVert_{\boldsymbol{C}_{i}^{\ast}}^{2}\right)$
where $\boldsymbol{C}^{\ast}_{i}=\operatorname{Exp}_{\boldsymbol{C}^{\ast}}(\sum_{k}x_{k}v_{k})$.
Existence and uniqueness of OLS estimates in general is known (Thomas Fletcher, 2013). Of course, with this likelihood they are MLE estimators.
We may relax the sphericity assumption by assumming instead that there is a positive definite $\Gamma$ such that
$p(\boldsymbol{C}_{i})\propto\exp\left(-\frac{1}{2}\operatorname{Vect}_{\boldsymbol{C}_{i}^{\ast}}(\log_{\boldsymbol{C}_{i}^{\ast}}(\boldsymbol{C}_{i}))^{T}\boldsymbol{\Gamma}^{-1}\operatorname{Vect}_{\boldsymbol{C}_{i}^{\ast}}(\log_{\boldsymbol{C}_{i}^{\ast}}(\boldsymbol{C}_{i}))\right)$ (1)
In this case, we have
Proposition 2.2.
There exist unique MLE estimators for model (1)
Proof.
Consider the affine invariant Riemannian metric $\langle,\rangle^{\operatorname{Mah}}$, defined on $\mathcal{M}$ by requiring
$\langle\log_{\mathbf{I}}(\boldsymbol{C}),\log_{\mathbf{I}}(\boldsymbol{\Lambda})\rangle_{\mathbf{I}}^{\operatorname{Mah}}=\operatorname{Vect}_{\mathbf{I}}(\log_{\mathbf{I}}(\boldsymbol{C}))^{T}\boldsymbol{\Gamma}^{-1}\operatorname{Vect}_{\mathbf{I}}(\log_{\mathbf{I}}(\boldsymbol{\Lambda}))$
Then the MLE estimators of model (1) are the OLS estimates of $\langle,\rangle^{\operatorname{Mah}}$, whose existence and uniqueness follow from (Thomas Fletcher, 2013). ∎
We call the family of metrics introduced in the previous proof Mahalanobis Affine Invariant Riemannian Metric (MAIRM) and the likelihood above the Riemannian normal model.
2.2 Riemannian Generalization of Wilk’s Lambda
Let us now extend the concepts above to develop a multivariate test statistic that generalizes the classical Wilk’s Lambda to the SPD manifold. This generalization enables us to test for differences among multiple groups of connectomes while respecting the natural geometric structure of the data.
To introduce the statistic that generalizes Wilk’s Lambda, let us set up some notation. In the manifold $\mathcal{M}=\operatorname{SPD}$, consider $\mathcal{M}$-valued random variables $C_{lj}$, $l=1,\ldots,g$, $j=1,\ldots,n_{l}$.
The usual orthogonal decomposition that is the basis of ANOVA does not directly apply in this manifold setting, and we need to find a suitable substitute.
Let $\hat{C}$ denote the Fréchet sample mean of the whole sample (see Appendix 1 for details) and $\hat{C}_{l}$ denote the Fréchet sample mean of the subsample $\{C_{lj}\}_{j}$.
Proposition 2.3.
[…]
$T$ $=\sum_{l=1}^{g}\sum_{j=1}^{n_{l}}w_{lj}w_{lj}^{T}$
$W$ $=\sum_{l=1}^{g}\sum_{j=1}^{n_{l}}u_{lj}u_{lj}^{T}$
If $T$ is positive definite, then the test statistic
$\Lambda^{\ast}=\frac{|W|}{|T|}$
ranges between 0 and 1.
Proof.
Since the vectors $\{u_{lj}\}_{j}$ are centered (i.e., $\sum_{j=1}^{n_{l}}u_{lj}=0$), we have
$\sum_{j=1}^{n_{l}}w_{lj}w_{lj}^{T}=\sum_{j=1}^{n_{l}}(v_{l}-u_{lj})(v_{l}-u_{lj})^{T}$
$=n_{l}v_{l}v_{l}^{T}+\sum_{j=1}^{n_{l}}u_{lj}u_{lj}^{T}$ (2)
Adding over $l$ yields the decomposition
$T=\sum_{l=1}^{g}n_{l}v_{l}v_{l}^{T}+W$ (3)
Both matrices are positive semidefinite and equation (3) shows $T\geq W$, hence $0\leq\Lambda^{\ast}\leq 1$. ∎
Remark. Notice the difference between $w_{lj}$ and the vector $\operatorname{Vec}_{\hat{C}}(\log_{\hat{C}}C_{lj})$. In general, they are not the same, but in the Euclidean case they coincide, so the subsequent developments reduce to the traditional MANOVA in that case.
We consider the null hypothesis $H_{0}:C_{1}^{*}=C_{2}^{*}=\cdots=C_{g}^{*}$, where $C_{l}^{*}$ denotes the population Fréchet mean of group $l$. This is equivalent to testing whether all groups share the same population mean in the manifold.
2.3 Asymptotic Distribution of the Test Statistic
We now present the asymptotic distribution of our Riemannian ANOVA test statistic $\Lambda^{*}$ under this model.
Theorem 2.4 (Asymptotic Distribution).
Let $C_{lj}\sim p(C)$ for $l=1,\ldots,g$ and $j=1,\ldots,n_{l}$, where
$p(C)\propto\exp\left(-\frac{1}{2}\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(C))^{T}\Gamma^{-1}\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(C))\right)$
is the Riemannian normal distribution on the manifold of symmetric positive definite matrices with population mean $C^{*}$ and covariance $\Gamma$ in the tangent space.
Under the null hypothesis $H_{0}:C^{*}_{1}=C^{*}_{2}=\cdots=C^{*}_{g}=C^{*}$ as $n=\sum_{l=1}^{g}n_{l}\to\infty$ with $n_{l}/n\to\pi_{l}\in(0,1)$:
$n\log\Lambda^{*}\xrightarrow{d}\chi^{2}_{d(g-1)}$
where $d=p(p+1)/2$ is the dimension of $\mathcal{M}$.
Proof.
The proof proceeds by establishing the asymptotic behavior of the components of $\Lambda^{*}$ and then deriving the limiting distribution. Recall from Proposition 2.3 that the test statistic is defined as:
$\Lambda^{*}=\frac{|W|}{|T|}$
where $T=\sum_{l=1}^{g}\sum_{j=1}^{n_{l}}w_{lj}w_{lj}^{T}$ and $W=\sum_{l=1}^{g}\sum_{j=1}^{n_{l}}u_{lj}u_{lj}^{T}$.
We begin by applying the central limit theorem for Fréchet means (Bhattacharya and Patrangenaru, 2005). Define the chart $\phi:\text{SPD}(p)\to\mathbb{R}^{d}$ by $\phi(C)=\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(C))$.
We first calculate the gradient $\Psi(u;\theta):=\text{grad}_{\theta}\rho_{\phi}^{2}(u,\theta)$, where $\rho_{\phi}^{2}(u,\theta)=\rho^{2}(\text{Exp}_{C^{*}}(u),\text{Exp}_{C^{*}}(\theta))$ and $\rho$ denotes the geodesic distance on the manifold. We have
$\Psi(u;\theta)=-2G_{C^{*}}(u-\theta)+O(\|\theta\|^{2})$
where $G_{C^{*}}$ is the matrix representation of the metric tensor at $C^{*}$ in vectorized coordinates, arising from the inner product $\langle V,W\rangle_{C^{*}}=\text{tr}((C^{*})^{-1}V(C^{*})^{-1}W)$.
Next, under $H_{0}$, with $u=\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(Y))\sim N(0,\Gamma)$, we have:
$\Sigma=\text{Cov}(\Psi(u;0))=\text{Cov}(-2G_{C^{*}}u)=4G_{C^{*}}\Gamma G_{C^{*}}^{T}$
We now compute the Hessian $\Lambda$. The Hessian of the squared geodesic distance at the minimum is:
$\Lambda_{ij}=E\left[\frac{\partial}{\partial\theta_{j}}\Psi_{i}(u;\theta)\Big|_{\theta=0}\right]=2[G_{C^{*}}]_{ij}$
That is, $\Lambda=2G_{C^{*}}$.
Applying Theorem 2.1 of Bhattacharya and Patrangenaru (2003), we obtain:
$\sqrt{n}(\phi(\hat{C})-\phi(C^{*}))\xrightarrow{d}N(0,\Lambda^{-1}\Sigma(\Lambda^{T})^{-1})$
Substituting our computed values:
$\Lambda^{-1}\Sigma(\Lambda^{T})^{-1}=(2G_{C^{*}})^{-1}\cdot 4G_{C^{*}}\Gamma G_{C^{*}}^{T}\cdot(2G_{C^{*}}^{T})^{-1}=\Gamma$
That is, the metric terms cancel, which establishes that:
$)\displaystyle\sqrt{n}\,\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(\hat{C}$)) $\xrightarrow{d}N(0,\Gamma)$ (4)
With this result at hand, we now turn to the asymptotic behavior of tangent vectors. For the between-group deviation vectors:
$)v_{l}=\operatorname{Vec}_{\hat{C}_{l}}(\log_{\hat{C}_{l}}(\hat{C}$))
we proceed as follows.
First, using Corollary A.2, with $A=\hat{C}_{l}$, $B=\hat{C}$, $C=C^{*}$, and noting that $d_{R}(\hat{C}_{l},C^{*})=O_{p}(n_{l}^{-1/2})$ and $d_{R}(\hat{C},C^{*})=O_{p}(n^{-1/2})$ from the convergence of Fréchet sample means, we obtain
$\log_{\hat{C}_{l}}(\hat{C})=\Pi_{C^{*}\to\hat{C}_{l}}[\log_{C^{*}}(\hat{C})-\log_{C^{*}}(\hat{C}_{l})]+O_{p}(n^{-1/2})$
By the central limit theorem for Fréchet means we have $\log_{C^{*}}(\hat{C})=O_{p}(n^{-1/2})$ and $\log_{C^{*}}(\hat{C}_{l})=O_{p}(n_{l}^{-1/2})$. Therefore, the difference becomes:
$\log_{C^{*}}(\hat{C})-\log_{C^{*}}(\hat{C}_{l})=O_{p}(n^{-1/2})$
Combining all these results:
$)\displaystyle v_{l}=\operatorname{Vec}_{\hat{C}_{l}}(\log_{\hat{C}_{l}}(\hat{C}$)) $=\operatorname{Vec}_{C^{*}}\Pi_{C^{*}\to\hat{C}_{l}}(\log_{C^{*}}(\hat{C})-\log_{C^{*}}(\hat{C}_{l}))+O_{p}(n^{-1/2})$
$=O_{P}(n^{-1/2})+O_{p}(n^{-1/2})=O_{P}(n^{-1/2})$
We now consider the observation vectors:
$u_{lj}=\operatorname{Vec}_{\hat{C}_{l}}(\log_{\hat{C}_{l}}(C_{lj}))$
To establish the stability under change of base point, we analyze how the vectorization changes from $\hat{C}_{l}$ to $C^{*}$. Using Corollary A.2, we obtain the expansion:
$\log_{\hat{C}_{l}}(C_{lj})=\Pi_{C^{*}\to\hat{C}_{l}}\left(\log_{C^{*}}(C_{lj})-\log_{C^{*}}(C_{l})\right)+O(n^{-1/2})$
Therefore, combining these estimates:
$u_{lj}=\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(C_{lj}))+O_{p}(n_{l}^{-1/2})$
Next, we examine the decomposition and matrix distributions. The decomposition $T=B+W$ holds where:
$B$ $=\sum_{l=1}^{g}n_{l}v_{l}v_{l}^{T}=O_{p}(1)$ (5)
$W$ $=\sum_{l=1}^{g}\sum_{j=1}^{n_{l}}u_{lj}u_{lj}^{T}$ (6)
Due to the asymptotic normality in the tangent space, we have an approximate Wishart distribution for $W$. More precisely, the argument above gives:
$u_{lj}=z_{lj}+\varepsilon_{lj}$
where $z_{lj}\sim N(0,\Gamma)$ exactly and $\varepsilon_{lj}=O_{p}(n_{l}^{-1/2})$ represents the error from using the estimated group mean. This leads to the decomposition:
$W=\sum_{l=1}^{g}\sum_{j=1}^{n_{l}}u_{lj}u_{lj}^{T}=Z+E+R$
where $Z\sim\mathcal{W}_{d}(n-g,\Gamma)$ (the degrees of freedom are due to the centering constraints $\sum_{j=1}^{n_{l}}u_{lj}=0$) is the exact Wishart component, $E=O_{p}(n_{l}^{-1/2})$ represents cross-terms between true and error components, and $R=O_{p}(n_{l}^{-1/4})$ is the quadratic error term.
The rest of the argument proceeds exactly as in Euclidean MANOVA (Anderson and Olkin, 1994). ∎
2.4 Comparison to Fréchet ANOVA
The Fréchet ANOVA test (Dubey and Müller, 2019) provides a distribution-free alternative for testing equality of population Fréchet means and variances in general metric spaces. We now establish the precise relationship between Fréchet ANOVA and our Riemannian ANOVA test under the Riemannian normal model.
The Fréchet ANOVA test statistic takes the form:
$T_{n}=\frac{nU_{n}}{\sum_{l=1}^{g}\frac{\lambda_{l,n}}{\hat{\sigma}_{l}^{2}}}+\frac{nF_{n}^{2}}{\sum_{l=1}^{g}\lambda_{l,n}^{2}\hat{\sigma}_{l}^{2}}$
[…]
Proposition 2.5 (Fréchet ANOVA Non-centrality Under Normality).
Under the Riemannian normal model with equal covariances and alternative hypothesis $H_{1}$, the Fréchet ANOVA test statistic follows asymptotically:
$T_{n}\xrightarrow{d}\chi^{2}_{g-1}(\delta_{\text{Fréchet}})$
where the non-centrality parameter is:
$\delta_{\text{Fréchet}}=\frac{n\left(\sum_{l=1}^{g}\lambda_{l}d^{2}(C_{l}^{*},C^{*})\right)^{2}}{\operatorname{tr}(\Gamma)\sum_{l=1}^{g}\lambda_{l}^{2}}$
with $C^{*}$ denoting the overall population Fréchet mean.
Proof.
Suppose $C_{lj}\sim p(C_{l}^{*})$ where the connectomes follow a Riemannian normal distribution on $\mathrm{SPD}(p)$ with population mean $C_{l}^{*}$ and covariance $\Gamma$ in the tangent space (as defined in Section 2.3).
Under the equal covariance assumption, the population Fréchet variance for each group is:
$V_{l}=\mathbb{E}[d^{2}(C_{lj},C_{l}^{*})]=\operatorname{tr}(\Gamma)$
for all $l=1,\ldots,g$, since $\log_{C_{l}^{*}}(C_{lj})\sim\mathcal{N}(0,\Gamma)$ in the vectorized tangent space.
The overall population Fréchet variance admits the decomposition:
$V_{p}=\sum_{l=1}^{g}\lambda_{l}\mathbb{E}[d^{2}(C_{lj},C^{*})]=\sum_{l=1}^{g}\lambda_{l}\left(V_{l}+d^{2}(C_{l}^{*},C^{*})\right),$
(similar to equation (2)). This yields:
$V_{p}=\operatorname{tr}(\Gamma)+\sum_{l=1}^{g}\lambda_{l}d^{2}(C_{l}^{*},C^{*}).$
Consequently, the scale component of the Fréchet ANOVA statistic vanishes:
$U=\sum_{j<l}\frac{\lambda_{j}\lambda_{l}}{\sigma_{j}^{2}\sigma_{l}^{2}}(V_{j}-V_{l})^{2}=0,$
since $V_{j}=V_{l}=\operatorname{tr}(\Gamma)$ for all $j,l$. And the location component at the population level simplifies to:
$F=V_{p}-\sum_{l=1}^{g}\lambda_{l}V_{l}=\sum_{l=1}^{g}\lambda_{l}d^{2}(C_{l}^{*},C^{*}),$
representing the weighted sum of squared Riemannian distances from population group means to the overall population mean.
Under contiguous alternatives with $C_{l}^{*}=\exp_{C^{*}}(n^{-1/2}H_{l})$ for symmetric matrices $H_{l}\in T_{C^{*}}\mathrm{SPD}(p)$, we have
$d^{2}(C_{l}^{*},C^{*})=\|n^{-1/2}H_{l}\|_{F}^{2}=n^{-1}\|H_{l}\|_{F}^{2}.$
Thus, the location component becomes
$F=n^{-1}\sum_{l=1}^{g}\lambda_{l}\|H_{l}\|_{F}^{2}.$
The asymptotic distribution of the test statistic $T_{n}=nF_{n}^{2}/\hat{\sigma}^{2}$ then follows from the chi-squared approximation, where $\hat{\sigma}^{2}\to\operatorname{tr}(\Gamma)$, yielding
$T_{n}\xrightarrow{d}\chi^{2}_{g-1}\left(\frac{n\left(\sum_{l=1}^{g}\lambda_{l}\|H_{l}\|_{F}^{2}\right)^{2}}{\operatorname{tr}(\Gamma)\sum_{l=1}^{g}\lambda_{l}^{2}}\right),$
which gives the stated non-centrality parameter. ∎
Proposition 2.6 (Non-centrality under Alternative Hypothesis).
Under the alternative hypothesis $H_{1}$ with population Fréchet means $C_{1}^{*},\ldots,C_{g}^{*}\in\mathrm{SPD}(p)$ where at least two differ, let $C^{*}$ denote the overall population Fréchet mean and define the population tangent vectors
$v_{l}^{*}=\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(C_{l}^{*}))$
for $l=1,\ldots,g$. Then, as $n\to\infty$ with $n_{l}/n\to\pi_{l}>0$:
  1. 1.
    The test statistic satisfies
    $n\log\Lambda^{*}\xrightarrow{d}\chi^{2}_{d(g-1)}(\delta)$
    where $d=p(p+1)/2$ and the non-centrality parameter is
    $\delta=\lim_{n\to\infty}n\cdot\mathrm{tr}(W^{*-1}B^{*})=\lim_{n\to\infty}n\sum_{l=1}^{g}\pi_{l}(v_{l}^{*})^{T}(W^{*})^{-1}v_{l}^{*}$
    with $B^{*}=\sum_{l=1}^{g}\pi_{l}v_{l}^{*}(v_{l}^{*})^{T}$ and $W^{*}$ the population within-group covariance matrix in the tangent space. This represents the sum of squared Mahalanobis distances of group means from the overall mean in the tangent space.
  2. 2.
    The power function at significance level $\alpha$ is
    $\beta_{n}=P\left(-n\log\Lambda^{*}>\chi^{2}_{d(g-1),1-\alpha}\mid H_{1}\right)\to 1$
    as $n\to\infty$.
Proof.
Let us analyze the between-group scatter under $H_{1}$. The between-group scatter matrix is
$B=\sum_{l=1}^{g}n_{l}\hat{v}_{l}\hat{v}_{l}^{T}$
Since $\hat{v}_{l}\xrightarrow{p}v_{l}^{*}\neq 0$ for at least some $l$ under $H_{1}$, we have
$\frac{B}{n}=\sum_{l=1}^{g}\frac{n_{l}}{n}\hat{v}_{l}\hat{v}_{l}^{T}\xrightarrow{p}\sum_{l=1}^{g}\pi_{l}v_{l}^{*}(v_{l}^{*})^{T}=B^{*}$
where $B^{*}$ is positive semi-definite with rank $\geq 1$. Thus $B=O_{p}(n)$.
For the within-group scatter, we have the matrix
$W=\sum_{l=1}^{g}\sum_{i=1}^{n_{l}}(v_{li}-\hat{v}_{l})(v_{li}-\hat{v}_{l})^{T}$
where $v_{li}=\operatorname{Vec}_{\hat{C}}(\log_{\hat{C}}(C_{li}))$ is the tangent vector representation of the $i$-th observation from group $l$ at the overall sample mean $\hat{C}$. This converges as $W/n\xrightarrow{p}W^{*}$ where
$W^{*}=\sum_{l=1}^{g}\pi_{l}C_{l}$
with $C_{l}=E[(v_{li}-v_{l}^{*})(v_{li}-v_{l}^{*})^{T}]$ being the population covariance within group $l$. Thus $W=O_{p}(n)$.
We now examine the limiting behavior of Wilks’ Lambda. Since both $W$ and $B$ are $O_{p}(n)$, we can write
$\Lambda^{*}=\frac{|W|}{|W+B|}=\frac{|\tilde{W}|}{|\tilde{W}+\tilde{B}|}$
where $\tilde{W}=W/n$ and $\tilde{B}=B/n$. As $n\to\infty$:
$\Lambda^{*}\xrightarrow{p}\frac{|W^{*}|}{|W^{*}+B^{*}|}=\lambda^{*}<1$
Since $B^{*}>0$ under $H_{1}$, we have $\lambda^{*}<1$ strictly.
This leads us to the asymptotic distribution. The log-likelihood ratio statistic
[…]
where the factor of $n$ arises because we express the determinants in terms of $\tilde{W}=W/n$ and $\tilde{B}=B/n$, which converge to finite limits. Using the expansion $\log|I+A|\approx\mathrm{tr}(A)-\frac{1}{2}\mathrm{tr}(A^{2})+O(\|A\|^{3})$ for the matrix $A=\tilde{W}^{-1}\tilde{B}$, we obtain the leading term $n\cdot\mathrm{tr}(W^{*-1}B^{*})$.
Finally, Since $\delta_{n}=O(n)$, the non-centrality parameter grows linearly with sample size. Therefore, for any fixed critical value $c_{\alpha}$:
$\beta_{n}=P(-n\log\Lambda^{*}>c_{\alpha}\mid H_{1})\to 1$
as $n\to\infty$. ∎
Proposition 2.7 (Comparison of Non-centrality Parameters).
Consider the Riemannian ANOVA and Fréchet ANOVA tests under the Riemannian normal model with isotropic covariance $\Gamma=\sigma^{2}I_{d^{2}}$ and balanced design ($n_{l}=n/g$ for all groups). For a specific alternative where only group 1 deviates from the common mean by Riemannian distance $\theta>0$ (i.e., $d(C_{1}^{*},C^{*})=\theta$ and $d(C_{l}^{*},C^{*})=0$ for $l=2,\ldots,g$), the non-centrality parameters simplify to:
$\delta_{\text{Novel}}$ $=\frac{n\theta^{2}}{g\sigma^{2}}=O(\theta^{2})$ (7)
$\delta_{\text{Fréchet}}$ $=\frac{n\theta^{4}}{g\sigma^{2}d^{2}}=O(\theta^{4})$ (8)
Proof.
We establish the non-centrality expressions for both tests under the specified conditions.
The general expression for the non-centrality parameter can be rewritten as:
$\delta_{\text{Novel}}=n\sum_{l=1}^{g}\lambda_{l}\|v_{l}^{*}\|_{\Gamma^{-1}}^{2}$
where $v_{l}^{*}=\operatorname{Vec}(\log_{C^{*}}(C_{l}^{*}))$ and $\lambda_{l}=1/g$ for balanced designs.
For the specific alternative with only group 1 deviating:
$\|v_{1}^{*}\|_{\Gamma^{-1}}^{2}=\frac{1}{\sigma^{2}}\|v_{1}^{*}\|_{F}^{2}=\frac{\theta^{2}}{\sigma^{2}}$
Therefore:
$\delta_{\text{Novel}}=\frac{n\theta^{2}}{g\sigma^{2}}$
On the other hand, the Fréchet ANOVA non-centrality parameter is:
$\delta_{\text{Fréchet}}=\frac{n\left(\sum_{l=1}^{g}\lambda_{l}d^{2}(C_{l}^{*},C^{*})\right)^{2}}{\operatorname{tr}(\Gamma)\sum_{l=1}^{g}\lambda_{l}^{2}}$
Substituting these values:
$\delta_{\text{Fréchet}}=\frac{n\cdot(\theta^{2}/g)^{2}}{\sigma^{2}d^{2}\cdot(1/g)}=\frac{n\theta^{4}}{g\sigma^{2}d^{2}}$
Remark 2.8 (Practical Implications).
The quadratic versus quartic dependence on $\theta$ has significant practical consequences. For detecting small location shifts, the Riemannian ANOVA test requires substantially smaller sample sizes to achieve the same power as Fréchet ANOVA. However, Fréchet ANOVA retains advantages in robustness to model misspecification and the ability to detect variance heterogeneity through its scale component $U_{n}$. The choice between tests should consider both the expected effect size and confidence in the Riemannian normality assumption.
3 Conclusion
In this paper, we have shown that while Fréchet ANOVA is a widely applicable test, we can sacrifice some of its generality to obtain a test that in certain situations can be more powerful. We have proposed one such test by leveraging properties of certain Riemannian metrics on the SPD manifold to carry over the sum of squares decomposition at the heart of classical multivariate ANOVA. We have also identified distributional assumptions, namely the Riemannian normal model, under which we can derive the asymptotic distribution of the novel test statistic. Substantial technical work was done in this direction, consisting of analyzing the behavior of certain Jacobi fields to bound the possible effect of curvature on the classical decomposition of observations from the Fréchet mean. In turn, we have used that asymptotic distribution to show that the novel test is more powerful than Fréchet ANOVA in certain regions of the parameter space. In this way, we have shown that the study of the statistical properties of manifold-valued samples still offers plenty of new research directions to explore, and that the geometry of the sample space endows statisticians with a varied and powerful toolkit.
Appendix A Riemannian Geometry Background
A.1 Riemannian Geometry Review
For a parcellation of size $p$, connectomes are symmetric, positive definite, $p\times p$ matrices. The set $\mathcal{M}=\mathrm{SPD}_{p}$ is a convex cone, but it is not a linear subspace of the set of $p\times p$ matrices, $\mathrm{Mat}\left(p\right)$. They are an open subset of the linear subspace of $p\times p$ symmetric matrices, $\mathrm{Symm}_{p}$. As such, they constitute a smooth manifold of dimension $d=\frac{p\left(p+1\right)}{2}$. We denote a generic point of $\mathcal{M}$ as $C$. The tangent space at any $C$, $T_{C}\mathcal{M}$, is naturally isomorphic to $\mathrm{Symm}_{p}$.
Furthermore, $\mathcal{M}$ can be endowed with the structure of a Riemannian manifold by choosing a Riemannian metric. As mentioned before, a popular choice is AIRM (Pennec, Fillard and Ayache, 2006), which can be introduced in a couple of ways. On one hand, it is a Fisher information metric related to the fact that each $C\in\mathcal{M}$ is the set of parameters for a distribution, namely $N\left(0,C\right)$ (Lenglet et al., 2006b). On the other hand, $\mathcal{M}$ is a homogeneous space of the Lie group $\mathrm{Mat}\left(p\right)^{\times}$ under the action given by $C\mapsto ACA^{T}$ for $A\in\mathrm{Mat}\left(p\right)^{\times}$. The left invariant metric on $\mathrm{Mat}\left(p\right)^{\times}$ descends to a metric on $\mathcal{M}$.
AIRM has a prominent role in and it has several nice properties. It is a complete metric: geodesics stay on the manifold for infinite time (the boundary of $\mathcal{M}$, the set of semidefinite matrices, are at infinite distance in AIRM). Also, it has non-positive curvature. These two facts allow us to use Hadamard’s theorem to conclude that for each $C$, the maps $\exp_{C}:\mathcal{M}\rightarrow T_{C}\mathcal{M}$ and $\log_{C}:T_{C}\mathcal{M}\rightarrow\mathcal{M}$ are global, mutually inverse diffeomorphisms. AIRM introduces a geodesic distance $d\left(\cdot,\cdot\right)$ on $\mathcal{M}$ as well as a norm $\left\lVert\cdot\right\rVert_{C}$ on $T_{C}\mathcal{M}$, which satisfy the relation
$\left\lVert\log_{C}\left(\Lambda\right)\right\rVert_{C}=d\left(C,\Lambda\right)$
AIRM allows us to introduce several probabilistic features which we can use to perform statistical inference on them. First, it allows us to compute expected values. We cannot perform integration due to the lack of a linear structure. Instead, we use the notion of Fréchet mean: for a random connectome $C$,
$\mathbb{E}\left[C\right]=\operatorname{argmin}_{\Lambda}\left\lVert\log_{\Lambda}\left(C\right)\right\rVert_{\Lambda}^{2}$
We can also introduce a notion of normal distribution, as follows. AIRM provides us with a linear isometry $\operatorname{Vec}_{C}:T_{C}\mathcal{M}\rightarrow\mathbb{R}^{p}$. For matrices $C^{\ast}\in\mathcal{M},\Gamma\in\mathrm{SPD}_{d}$, we say the random connectome $C$ has distribution $N\left(C^{\ast},\Gamma\right)$ if the random vector $\operatorname{Vec}_{C^{\ast}}\left(\log_{C^{\ast}}\left(C\right)\right)$ has distribution $N\left(0,\Gamma\right)$.
A.2 Technical Results
The proof of Theorem 2.4 requires the following estimates.
Theorem A.1 (Jacobi Field Quotient Identity).
Let $M$ be a Riemannian manifold and $\gamma:[0,1]\to M$ be the geodesic $\gamma(s)=\exp_{C}(s\log_{C}(A))$ connecting $C$ to $A$. Consider the variation $\sigma(s,t)=\exp_{\gamma(s)}(t\log(B))$ Let $J(s)=\frac{\partial\sigma(s,t)}{\partial t}\big|_{t=0}$ be the associated Jacobi field. Then for small $\varepsilon>0$:
$\frac{\log_{\gamma(\varepsilon)}(B)-\Pi_{C\to\gamma(\varepsilon)}\left[\log_{C}(B)\right]}{\varepsilon}=\frac{D}{ds}J\bigg|_{s=\varepsilon}+O(\varepsilon)$
where $\Pi_{C\to\gamma(\varepsilon)}$ denotes parallel transport along $\gamma$ from $C$ to $\gamma(\varepsilon)$.
Proof.
We establish the result through a careful analysis of the Jacobi field evolution and its covariant derivatives.
Let $\{E_{i}(s)\}_{i=1}^{n}$ be a parallel orthonormal frame along $\gamma$, constructed by parallel transporting an orthonormal basis from $T_{C}M$ along the geodesic. Since the frame is parallel, we have $\frac{D}{ds}E_{i}=0$ for all $i$ and $s\in[0,1]$.
Express the Jacobi field in this frame:
$J(s)=\sum_{i=1}^{n}J^{i}(s)E_{i}(s)$
where $J^{i}(s)=\langle J(s),E_{i}(s)\rangle$ are the component functions.
Since the frame is parallel, the covariant derivatives of $J$ have a simple form:
$\frac{D}{ds}J=\sum_{i=1}^{n}\frac{dJ^{i}}{ds}E_{i}(s),\quad\frac{D^{2}}{ds^{2}}J=\sum_{i=1}^{n}\frac{d^{2}J^{i}}{ds^{2}}E_{i}(s)$
The Jacobi equation $\frac{D^{2}}{ds^{2}}J+R(J,\dot{\gamma})\dot{\gamma}=0$ becomes:
$\sum_{i=1}^{n}\frac{d^{2}J^{i}}{ds^{2}}E_{i}(s)+R(J(s),\dot{\gamma}(s))\dot{\gamma}(s)=0$
In the parallel frame, the curvature endomorphism $R(\cdot,\dot{\gamma})\dot{\gamma}$ has components:
$R(J,\dot{\gamma})\dot{\gamma}=\sum_{i,j=1}^{n}R_{ij}J^{j}E_{i}$
where $R_{ij}=\langle R(E_{j},\dot{\gamma})\dot{\gamma},E_{i}\rangle$.
Therefore, the component equations are:
$\frac{d^{2}J^{i}}{ds^{2}}+\sum_{j=1}^{n}R_{ij}(s)J^{j}(s)=0,\quad i=1,\ldots,n$
For each component function $J^{i}(s)$, we have the Taylor expansion around $s=0$:
$J^{i}(s)=J^{i}(0)+s\frac{dJ^{i}}{ds}\bigg|_{s=0}+\frac{s^{2}}{2}\frac{d^{2}J^{i}}{ds^{2}}\bigg|_{s=0}+O(s^{3})$
From the Jacobi equation at $s=0$:
$\frac{d^{2}J^{i}}{ds^{2}}\bigg|_{s=0}=-\sum_{j=1}^{n}R_{ij}(0)J^{j}(0)$
Substituting:
$J^{i}(\varepsilon)=J^{i}(0)+\varepsilon\frac{dJ^{i}}{ds}\bigg|_{s=0}-\frac{\varepsilon^{2}}{2}\sum_{j=1}^{n}R_{ij}(0)J^{j}(0)+O(\varepsilon^{3})$
The parallel transport of $J(0)\in T_{C}M$ to $T_{\gamma(\varepsilon)}M$ is given by:
$\Pi_{C\to\gamma(\varepsilon)}[J(0)]=\sum_{i=1}^{n}J^{i}(0)E_{i}(\varepsilon)$
This vector in $T_{\gamma(\varepsilon)}M$ has components $J^{i}(0)$ in the parallel frame at $\gamma(\varepsilon)$.
The Jacobi field at $\gamma(\varepsilon)$ is:
$J(\varepsilon)=\sum_{i=1}^{n}J^{i}(\varepsilon)E_{i}(\varepsilon)$
Both $J(\varepsilon)$ and $\Pi_{C\to\gamma(\varepsilon)}[J(0)]$ are vectors in $T_{\gamma(\varepsilon)}M$. Their difference is:
$J(\varepsilon)-\Pi_{C\to\gamma(\varepsilon)}[J(0)]=\sum_{i=1}^{n}[J^{i}(\varepsilon)-J^{i}(0)]E_{i}(\varepsilon)$
Substituting the expansion:
$J(\varepsilon)-\Pi_{C\to\gamma(\varepsilon)}[J(0)]=\varepsilon\sum_{i=1}^{n}\frac{dJ^{i}}{ds}\bigg|_{s=0}E_{i}(\varepsilon)-\frac{\varepsilon^{2}}{2}\sum_{i,j=1}^{n}R_{ij}(0)J^{j}(0)E_{i}(\varepsilon)+O(\varepsilon^{3})$
Note that $\sum_{i=1}^{n}\frac{dJ^{i}}{ds}\big|_{s=0}E_{i}(\varepsilon)$ is the parallel transport of $\frac{D}{ds}J\big|_{s=0}$ to $T_{\gamma(\varepsilon)}M$.
Dividing by $\varepsilon$:
$\frac{J(\varepsilon)-\Pi_{C\to\gamma(\varepsilon)}[J(0)]}{\varepsilon}=\sum_{i=1}^{n}\frac{dJ^{i}}{ds}\bigg|_{s=0}E_{i}(\varepsilon)-\frac{\varepsilon}{2}\sum_{i,j=1}^{n}R_{ij}(0)J^{j}(0)E_{i}(\varepsilon)+O(\varepsilon^{2})$
The covariant derivative $\frac{D}{ds}J$ along the geodesic has components $\frac{D}{ds}J]^{i}(s)=\frac{dJ^{i}}{ds}(s)$.
From the Jacobi equation:
$\frac{d}{ds}\left[\frac{dJ^{i}}{ds}\right]=\frac{d^{2}J^{i}}{ds^{2}}=-\sum_{j=1}^{n}R_{ij}(s)J^{j}(s)$
For small $\varepsilon$:
$\frac{D}{ds}J]^{i}(\varepsilon)=[\frac{D}{ds}J]^{i}(0)-\varepsilon\sum_{j=1}^{n}R_{ij}(0)J^{j}(0)+O(\varepsilon^{2})$
Therefore, $\frac{D}{ds}J\big|_{s=\varepsilon}$ expressed in the frame at $\gamma(\varepsilon)$ is:
$\frac{D}{ds}J\big|_{s=\varepsilon}=\Pi_{C\to\gamma(\varepsilon)}[\frac{D}{ds}J\big|_{s=0}]-\varepsilon\Pi_{C\to\gamma(\varepsilon)}[R(J(0),\dot{\gamma}(0))\dot{\gamma}(0)]+O(\varepsilon^{2})$
From the quotient expression and the evolution of $\frac{D}{ds}J$:
$\Pi_{C\to\gamma(\varepsilon)}[\frac{D}{ds}J\big|_{s=0}]=\frac{D}{ds}J\big|_{s=\varepsilon}+\varepsilon\Pi_{C\to\gamma(\varepsilon)}[R(J(0),\dot{\gamma}(0))\dot{\gamma}(0)]+O(\varepsilon^{2})$
Substituting:
$\frac{J(\varepsilon)-\Pi_{C\to\gamma(\varepsilon)}[J(0)]}{\varepsilon}=\frac{D}{ds}J\big|_{s=\varepsilon}+\frac{\varepsilon}{2}\Pi_{C\to\gamma(\varepsilon)}[R(J(0),\dot{\gamma}(0))\dot{\gamma}(0)]+O(\varepsilon^{2})$
The curvature term $\frac{\varepsilon}{2}\Pi_{C\to\gamma(\varepsilon)}[R(J(0),\dot{\gamma}(0))\dot{\gamma}(0)]$ is of order $\varepsilon$, so:
$\frac{J(\varepsilon)-\Pi_{C\to\gamma(\varepsilon)}[J(0)]}{\varepsilon}=\frac{D}{ds}J\big|_{s=\varepsilon}+O(\varepsilon)$
Finally, since $\gamma(\varepsilon)=\exp_{C}(\varepsilon\log_{C}(A))$ and $J(\varepsilon)=\log_{\gamma(\varepsilon)}(B)$ and $J(0)=\log_{C}(B)$:
$\frac{\log_{\gamma(\varepsilon)}(B)-\Pi_{C\to\gamma(\varepsilon)}[\log_{C}(B)]}{\varepsilon}=\frac{D}{ds}J\bigg|_{s=\varepsilon}+O(\varepsilon)$
Corollary A.2 (Logarithm Difference Formula).
Let $C,A,B\in M$ and consider the geodesic $\gamma(s)=\exp_{C}(s\log_{C}(A))$. For small $\varepsilon>0$, let $\gamma(\varepsilon)$ be a point on this geodesic close to $C$. Then:
$\frac{\log_{\gamma(\varepsilon)}(B)-\Pi_{C\to\gamma(\varepsilon)}[\log_{C}(B)-\log_{C}(A)]-\log_{\gamma(\varepsilon)}(A)}{\varepsilon}=\frac{D}{ds}\tilde{J}\bigg|_{s=\varepsilon}+O(\varepsilon)$
where $\tilde{J}(s)$ is the Jacobi field along $\gamma$ associated to the variation $\sigma(s,t)$ defined by:
$\sigma(s,t)=\exp_{\gamma(s)}(t[\log_{\gamma(s)}(B)-\log_{\gamma(s)}(A)])$
Proof.
This variation satisfies: $\sigma(s,0)=\gamma(s)$ for all $s$, $\sigma(0,t)=\exp_{C}(t[\log_{C}(B)-\log_{C}(A)])$, and $\sigma(1,t)=\exp_{A}(t\log_{A}(B))$ for all $t$ (since $\gamma(1)=A$ and $\log_{A}(A)=0$).
The associated Jacobi field is:
$\tilde{J}(s)=\frac{\partial\sigma}{\partial t}\bigg|_{t=0}=\log_{\gamma(s)}(B)-\log_{\gamma(s)}(A)$
Now we can decompose:
$\log_{\gamma(\varepsilon)}(B)=[\log_{\gamma(\varepsilon)}(B)-\log_{\gamma(\varepsilon)}(A)]+\log_{\gamma(\varepsilon)}(A)$
The first term is $\tilde{J}(\varepsilon)$, which by Theorem A.1 equals:
$\tilde{J}(\varepsilon)=\Pi_{C\to\gamma(\varepsilon)}[\tilde{J}(0)]+\varepsilon\frac{D}{ds}\tilde{J}\big|_{s=\varepsilon}+O(\varepsilon^{2})$
Since $\tilde{J}(0)=\log_{C}(B)-\log_{C}(A)$:
$\log_{\gamma(\varepsilon)}(B)-\log_{\gamma(\varepsilon)}(A)=\Pi_{C\to\gamma(\varepsilon)}[\log_{C}(B)-\log_{C}(A)]+\varepsilon\frac{D}{ds}\tilde{J}\big|_{s=\varepsilon}+O(\varepsilon^{2})$
Rearranging and dividing by $\varepsilon$:
$\frac{\log_{\gamma(\varepsilon)}(B)-\Pi_{C\to\gamma(\varepsilon)}[\log_{C}(B)-\log_{C}(A)]-\log_{\gamma(\varepsilon)}(A)}{\varepsilon}=\frac{D}{ds}\tilde{J}\bigg|_{s=\varepsilon}+O(\varepsilon)$
Proposition A.3 (AIRM Jacobi Field Velocity).
Let $(\mathcal{M},g^{\mathrm{AIRM}})$ be the manifold of $n\times n$ symmetric positive definite matrices equipped with the Affine Invariant Riemannian Metric (AIRM). Let $C,A,B\in\mathcal{M}$ be three points, and let $\gamma:[0,1]\to\mathcal{M}$ be the geodesic from $C$ to $A$, given by:
$\gamma(s)=C^{1/2}(C^{-1/2}AC^{-1/2})^{s}C^{1/2}$
Define the Jacobi field $J(s)$ along $\gamma$ arising from the variation:
$\sigma(s,t)=\exp_{\gamma(s)}(t[\log_{\gamma(s)}(B)-\log_{\gamma(s)}(A)])$
Then
$\frac{D}{ds}J\bigg|_{s=0}=-\frac{1}{2}[\log_{C}(A),\log_{C}(B)]-\frac{1}{2}\{\log_{C}(A),\log_{C}(B)\}_{C}+\log_{C}(A)$
where:
  • [X,Y]=XYYX[X,Y]=XY-YX is the matrix commutator
  • $\{X,Y\}_{C}=C^{-1}XY+YX^{T}C^{-1}$ is the symmetrized product with the metric at $C$
  • $\log_{P}(Q)=P^{1/2}\log(P^{-1/2}QP^{-1/2})P^{1/2}$ is the Riemannian logarithm map
Proof.
We denote $L_{A}=\log(C^{-1/2}AC^{-1/2})$ as the matrix logarithm, $L_{B}=\log(C^{-1/2}BC^{-1/2})$, and $M=C^{-1/2}AC^{-1/2}$ as the conjugated target matrix. Then $\log_{C}(A)=C^{1/2}L_{A}C^{1/2}$, $\log_{C}(B)=C^{1/2}L_{B}C^{1/2}$, $\gamma(s)=C^{1/2}M^{s}C^{1/2}$, and $\dot{\gamma}(s)=C^{1/2}L_{A}M^{s}C^{1/2}$.
From the definition of $\sigma$,
$J(s)=\log_{\gamma(s)}(B)-\log_{\gamma(s)}(A)$
For the AIRM metric at point $C\in\mathcal{M}$, the Levi-Civita connection is:
$\nabla_{X}Y=DY[X]-\frac{1}{2}(C^{-1}XY+YX^{T}C^{-1})$
where $X,Y\in T_{C}\mathcal{M}$ are tangent vectors (symmetric matrices) and $DY[X]$ is the Euclidean directional derivative.
Then
$\frac{D}{ds}\bigg|_{s=0}\log_{\gamma(s)}(B)=-\frac{1}{2}(C^{-1}\log_{C}(A)\log_{C}(B)+\log_{C}(B)\log_{C}(A)^{T}C^{-1})$
Decomposing the product as:
$\log_{C}(A)\log_{C}(B)=\frac{1}{2}[\log_{C}(A),\log_{C}(B)]+\frac{1}{2}\{\log_{C}(A),\log_{C}(B)\}$
where $\{X,Y\}=XY+YX$ is the anticommutator. We obtain:
$\frac{D}{ds}\bigg|_{s=0}\log_{\gamma(s)}(B)=-\frac{1}{2}[\log_{C}(A),\log_{C}(B)]-\frac{1}{2}\{\log_{C}(A),\log_{C}(B)\}_{C}$
Similarly,
$\frac{D}{ds}\bigg|_{s=0}\log_{\gamma(s)}(A)=-\log_{C}(A)$
And the result follows by gathering the terms.
Remark A.4.
The formula in Proposition A.3 is explicit. All terms can be computed using standard matrix operations: matrix logarithms (computable via eigendecomposition), matrix multiplications, and additions. No differential equation solving is required, making the computation both efficient and numerically stable.
{funding}
This research was supported in part by NIH grant R01 NS112303.